### v1

# Does Your Multi-modal LLM Truly See the Diagrams in Visual Math Problems?

Abstract
The remarkable progress of Multi-modal Large Language Models (MLLMs) has garnered unparalleled attention, due to their superior performance in visual contexts. However, their capabilities in visual math problem-solving remain insufficiently evaluated and understood. We investigate current benchmarks to incorporate excessive visual content within textual questions, which potentially assist MLLMs in deducing answers without truly interpreting the input diagrams. To this end, we introduce MathVerse, an all-around visual math benchmark designed for an equitable and in-depth evaluation of MLLMs. We meticulously collect 2,612 high-quality, multi-subject math problems with diagrams from publicly available sources. Each problem is then transformed by human annotators into six distinct versions, each offering varying degrees of information content in multi-modality, contributing to 15K test samples in total. This approach allows MathVerse to comprehensively assess whether and how much MLLMs can truly understand the visual diagrams for mathematical reasoning. In addition, we propose a Chain-of-Thought (CoT) evaluation strategy for a fine-grained assessment of the output answers. Rather than naively judging True or False, we employ GPT-4(V) to adaptively extract crucial reasoning steps, and then score each step with detailed error analysis, which can reveal the intermediate CoT reasoning quality by MLLMs. With MathVerse, we unveil that, most existing MLLMs struggle to understand math diagrams, relying heavily on textual questions. Surprisingly, some of them even achieve 5%+ higher accuracy without the visual input, e.g., Qwen-VL-Max and InternLM-XComposer2. In contrast, GPT-4V and ShareGPT4V demonstrate relatively better comprehension of the visual content for mathematical reasoning. We hope MathVerse may provide unique insights to guide the future development of MLLMs. Project page: https://mathverse-cuhk.github.io.
1 Introduction
With the substantial advances of big data and computational power, Large Language Models (LLMs) brown2020language ; albert24mixtral ; touvron2023llama ; touvron2023llama2 ; vicuna2023 , such as ChatGPT OpenAI2023ChatGPT and GPT-4 OpenAI2023GPT4TR , have emerged as a central point of interest in both industry and academia. To broaden their applicability across diverse contexts, Multi-modal Large Language Models (MLLMs) zhang2024llamaadapter ; gao2023llamaadapterv2 ; su2023pandagpt ; chen2023minigpt ; ye2023mplugowl have recently become a fast-evolving track, exemplified by the latest GPT-4V openai2023gpt4v , Gemini team2023gemini , and the open-source LLaVA liu2023llava ; liu2023improvedllava ; liu2024llavanext and SPHINX lin2023sphinx ; gao2024sphinx . Concurrently, a diverse array of evaluation benchmarks fu2023mme ; liu2023mmbench ; Li2023SEEDBenchBM ; fu2023challenger ; sun2024journeydb are curated to assess their visual comprehension performance across different domains. Notably, the capability to solve mathematical problems involving diagrams serves as a critical measure, offering insights into the multi-modal logical thinking prowess of MLLMs. This task demands MLLMs to accurately decode the visual elements within input diagrams (characters and figures), and correlate them with the condition specified by textual questions for mathematical reasoning. Previous efforts lu2021inter ; seo2015solving , e.g., GeoQA chen2021geoqa ; cao2022augmented and UniGeo chen2022unigeo , concentrate on the challenging geometric problems, while the recent MathVista Lu2023MathVistaEM and MMMU yue2023mmmu expand the scope to encompass broader disciplines, including functions, charts, and scientific problems.
Figure 1: (a) We showcase three examples of Text Redundancy (highlighted in red) within existing visual math benchmarks chen2021geoqa ; Lu2023MathVistaEM ; yue2023mmmu . (b) We report an ablation study by respectively removing the redundant texts and input diagrams on 120 randomly selected problems, for closed-sourced openai2023gpt4v ; team2023gemini ; bai2023qwen and open-sourced gao2024sphinx ; liu2024llavanext ; dong2024internlm MLLMs.
However, through our comprehensive observation and analysis, we identify three primary issues in current mathematical benchmarks for evaluating MLLMs:
  1. i.
    Do MLLMs truly see the math diagrams in evaluation? This is the most fundamental question concerning the accurate assessment of visual math problem-solving. In Figure 1 (a), we showcase three examples from current benchmarks. We observe their texts contain too much duplicate information (highlighted in red) that is also depicted in the diagram. This redundancy might inadvertently provide MLLMs with a shortcut to resolve the problem by mostly reading the text, rather than interpreting the diagram. Our hypothesis gains support from the experiment in Figure 1 (b). For 40 randomly sampled problems from each benchmark, we remove such redundant texts from the question, challenging MLLMs to capture the corresponding information exclusively from visual inputs. The results reveal a significant drop in accuracy among most MLLMs (the blue column), even falling below the scores without taking diagrams as input (the grey column). This outcome suggests that MLLMs primarily depend on textual cues rather than the visual diagrams themselves to solve these problems in evaluation. Given this, we demonstrate that current visual math benchmarks might not be comprehensive enough to assess the genuine multi-modal mathematical reasoning capabilities of MLLMs.
  2. ii.
    Is it equitable to assess solely by the final answer? Most existing multi-modal benchmarks directly compare model outputs with ground truths to derive a binary evaluation result. While this approach may suffice for general visual contexts, it falls short in math problems that require intricate step-by-step reasoning. In Figure 2, we examine three model outputs. Although they all arrive at incorrect answers in the end, they demonstrate varying levels of precision in the intermediate reasoning processes. Merely categorizing these outputs as ‘Incorrect’ fails to capture the nuanced differences in the reasoning quality of MLLMs.
  3. iii.
    Do they specialize in mathematical reasoning evaluation? GeoQA, UniGeo, and other previous works narrowly target specific aspects of plane geometry. This limits the evaluation of broader mathematical capabilities, e.g., functions and solid geometry. Instead, MathVista expands its scope by including a wide array of peripheral tasks (19 out of 28), encompassing natural images, statistic plots, and charts, which do not directly evaluate professional math skills. Furthermore, the math problems in MMMU are of college-level complexity with extensive domain-specific knowledge, potentially hindering MLLMs from fully demonstrating their reasoning capacity.
Figure 2: Comparison of Visual Mathematical Reasoning by Three MLLMs. Despite the incorrect final answer, GPT-4V openai2023gpt4v , Gemini-Pro team2023gemini , and SPHINX-MoE gao2024sphinx exhibit different levels of quality in the intermediate reasoning process.
Therefore, in light of the issues discussed, we present MathVerse, a holistic and specialized visual math benchmark crafted to evaluate the multi-modal mathematical reasoning skills of MLLMs. This benchmark encompasses a meticulously collected dataset of 2,612 visual math problems, with 1,236 newly acquired from public question repositories and 1,376 selected from existing benchmarks, ensuring a diverse range of challenges. To specialize in mathematical reasoning, MathVerse spans three primary areas: plane geometry, solid geometry, and functions. Each problem has been rigorously reviewed by expert annotators and classified into twelve detailed categories, emphasizing different fine-grained problem-solving capabilities. Notably, MathVerse distinguishes itself by introducing two novel strategies for evaluating MLLMs.
First, we investigate the influence of textual redundancy and validate whether MLLMs can interpret the diagrams for mathematical reasoning. As illustrated in Figure 3 (Left), we categorize the textual content within the questions into three different types: Descriptive Information, Implicit Property, and Essential Condition. These categories, arranged in ascending order of significance for problem-solving, correspond to information directly observable from the diagram, implicit spatial properties that demand advanced visual perception, and specific measurements crucial for computing the solution, respectively. Based on this problem formulation, expert annotators progressively remove the textual information from the questions in MathVerse, while incrementally incorporating elements into the visual diagrams to ensure problems are adequately defined. As shown in Figure 3 (Right), this process results in six unique versions of each problem characterized by a reduction in textual content and an enhancement in visual elements, creating a total of 15K test samples. These delicately curated problems can indicate the various multi-modal capabilities of MLLMs, such as geometric element understanding, function curve perception, and numerical value recognition, which thoroughly unveils whether and how much they comprehend the visual diagram for mathematical reasoning.
Second, to rigorously assess the visual Chain-of-Thought (CoT) capabilities wei2022chain , we propose a CoT Evaluation strategy for the step-by-step reasoning assessment of MLLMs. For each model’s output, we leverage GPT-4 to first extract several crucial steps exclusively from the solving process, deliberately omitting the input of the question and answer. This approach aims to mitigate the bias towards GPT-4’s inherent question-answering propensities. Then, the corresponding question, diagram, and ground-truth answer are fed into GPT-4 to evaluate each identified critical step, and provide detailed error analysis. Finally, the overall score is obtained by considering every single step within reasoning. Note that, we do not pre-define a ground-truth key-step template, since each math problem may encompass a variety of solution pathways, and different MLLMs tend to exhibit variable reasoning lengths. With CoT scoring, MathVerse showcases a fine-grained evaluation of the intermediate logical deduction of MLLMs, demonstrating visual mathematical CoT capabilities.
Figure 3: Three Categories of Question Texts in MathVerse. According to the significance for problem-solving, we categorize the question texts into three categories, and transform each problem into six versions for evaluation, with varying content in multi-modality. We present three examples in MathVerse for illustration.
We conduct extensive experiments on MathVerse with popular closed-source openai2023gpt4v ; bai2023qwen ; team2023gemini and open-source liu2023improvedllava ; liu2024llavanext ; dong2024internlm ; gao2024sphinx MLLMs. Comparing different problem versions, we unveil that, most existing MLLMs struggle to understand math diagrams, relying heavily on textual questions. Therein, GPT-4V openai2023gpt4v achieves the best overall performance across different problem versions and subjects. Surprisingly, some of the MLLMs even attain much higher results without the diagram input, e.g., +5.1% for Qwen-VL-Max bai2023qwen and +5.6% for InternLM-XComposer2 dong2024internlm . With the fine-grained error analysis produced by our CoT evaluation strategy, we demonstrate such results are due to their deficient visual encoding capacity for mathematical diagrams, which instead acts as a distraction for problem-solving. In contrast, GPT-4V and ShareGPT4V Chen2023ShareGPT4VIL demonstrate relatively better comprehension of the visual content for mathematical reasoning. Our experimental results suggest that inadequate mathematical visual interpretation capabilities represent the most significant impediment for MLLMs in addressing multi-modal math problems, indicating substantial potential for advancement.
The contributions of this paper are summarized as follows:
  • We investigate primary issues within existing benchmarks and introduce MathVerse, an all-around multi-modal benchmark evaluating the visual mathematical reasoning of MLLMs. The meticulously curated dataset contains 20K test problems with diagrams for a comprehensive assessment.
  • By modifying problems with varying information content in multi-modality, we explore whether and how much MLLMs can understand the visual diagrams for mathematical reasoning, rather than relying on question texts.
  • We propose a CoT evaluation strategy with GPT-4 to extract and assess each key step in the reasoning process of MLLMs, which provides a detailed error analysis and fine-grained evaluation of their multi-modal mathematical CoT capabilities.
2 MathVerse
In Section 2.1, we first present an overview of the curated visual math dataset in MathVerse. Then, in Section 2.2, we introduce our data formulation approach for investigating the visual mathematical comprehension of Multi-modal Large Language Models (MLLMs). Finally, in Section 2.3, we elaborate on the methodology of our proposed Chain-of-Thought (CoT) evaluation strategy.
[…]
2.1 Visual Math Dataset
To thoroughly assess visual mathematical proficiency, we compile a comprehensive problem set covering a broad spectrum of math subjects, diagram patterns, and specialized knowledge domains. This widespread collection for MathVerse aims to pose diverse challenges to MLLMs, ensuring a robust evaluation of their capabilities in visual contexts.
Table 1: Key Statistics of MathVerse.
Figure 4: Subject Distribution of MathVerse. Solid G: Solid Geometry, Plane G: Plane Geometry.
Statistic Number Total questions 2,612  - Multiple-choice questions 1,631 (62.4%)  - Free-form questions 981 (37.6%)  - Newly collected questions 1,236 (47.3%)  - Existing-dataset questions 1,376 (52.7%)  - Questions with explanations 1,236 (47.3%) Total test samples 15,672  - Newly annotated samples 10,448 (66.7%)  - Samples of each version 2,612 (16.7%) Number of unique images 2,420 (92.6%) Number of unique questions 2,573 (98.5%) Number of unique answers 847 (32.4%) Maximum question length 203 Maximum answer length 17 Average question length 35.7 Average answer length 1.4
Figure 4: Subject Distribution of MathVerse. Solid G: Solid Geometry, Plane G: Plane Geometry.
Data Composition and Categorization.
MathVerse comprises a total of 2,612 visual math problems, which contribute to the final created 15K test samples. Detailed statistics for data composition are presented in Table 4. This meticulously collected dataset covers three fundamental math subjects, i.e., plane geometry (1,746), solid geometry (332), and functions (534), where the latter two are all composed of newly collected problems. The choice of these three subjects is not only due to their rigorous demands on multi-modal reasoning, but also for two other considerations. For one thing, as we specialize MathVerse in mathematical problem-solving, other peripheral tasks in MathVista Lu2023MathVistaEM are not included, e.g., statistical reasoning, table question-answering, and puzzle tests. For another, we expect the evaluation to fully display the reasoning capabilities of MLLMs with moderate-level mathematical knowledge. This avoids limiting their performance with overly complex domain-specific theorems or prior commonsense knowledge. Therefore, we deliberately focus the collected problems on the high school level, excluding advanced college-level disciplines like calculus and graph theory featured in MMMU yue2023mmmu . Furthermore, expert annotators subdivide the problems into twelve fine-grained categories, as depicted in Figure 4, showcasing various dimensions of visual mathematical skills.
Data Collection and Review Process.
[…]
2.2 Whether MLLMs Truly See the Diagrams?
In this section, we detail our data formulation approach to transform each problem in MathVerse into six different versions with varying information content in multi-modality. In this way, we explore the visual diagram understanding capabilities of MLLMs for mathematical reasoning.
Figure 5: Six Versions of Each Problem in MathVerse. Expert annotators meticulously transform each visual math problem within MathVerse into six versions. They contain different vision-language content for a holistic visual mathematical evaluation.
Three Types of Textual Information.
Considering the textual redundancy in original math problems, we first define three distinct categories for the textual information within the questions, as illustrated in Figure 3 and the following:
  • Descriptive Information (DI) refers to the directly observable and clearly portrayed content in the diagram. It depicts the basic figure composition, spatial arrangement, and annotated entities, such as the presence of geometric shapes or intersection points of functions. These sentences normally help establish the context and frame the problem to orient the solver. Nevertheless, such information is repetitive to the visual components present in the diagram, thus regarded as redundant information for problem-solving. More importantly, it may assist MLLMs in bypassing the process of diagram interpretation, thereby undermining the assessment for visual mathematical reasoning, as evidenced in Figure 1.
  • Implicit Property (IP) involves the information that requires a higher level of visual perception but less mathematical knowledge to discern from the diagram. It signifies strong visual conditions for problem-solving, such as the parallelism and perpendicularity between lines, the similarity and congruence among triangles, and the category and periodicity of functions. They can, in theory, be fully extracted from the diagrams alone, giving adequate capability for visual recognition and comprehension of MLLMs.
  • Essential Condition (EC) denotes the specific numerical or algebraic measurements, which are indispensable conditions to derive the solution and cannot be derived from the visual diagram. This category encompasses precise values of angles, lengths, and function expressions, such as an angle being 45 degrees, the length of BC being 6 units, and the functional equation $f(x)=x^{2}+3$. Without these details in textual information, solving the visual math problem would be impossible.
Creating Six Versions of Each Problem.
Based on the three categories, expert annotators systematically remove different textual information within questions, and incrementally incorporate the critical elements into diagrams. This approach can progressively reduce textual redundancy and information content, thereby increasingly compelling MLLMs to capture mathematical conditions from the visual input. As compared in Figure 5, we generate six versions of each problem in MathVerse, obtaining 15,672 test instances. With this curated problem set, we can provide a holistic evaluation of the genuine visual comprehension of MLLMs, and whether it can facilitate multi-modal mathematical reasoning. The details of each problem version are as follows:
  • Text-dominant Version retains the entire textual content, including the three types of textual information and the question statement. If the original problem contains limited Descriptive Information, we manually add it within the textual content. This version may induce MLLMs to regard the text as the primary source of information, treating the diagram more as a supplementary visual aid. This serves as the baseline point for evaluation.
                        [Uncaptioned image] Text: DI + IP + EC + Question [Uncaptioned image] Vision: Diagram (1)
  • Text-lite Version diminishes the Descriptive Information from the Text-dominant version, assuming this information can be observed from the diagram. This creates a condensed question without redundancy, forcing MLLMs to interpret the diagram for basic information.
                        [Uncaptioned image] Text: IP + EC + Question        [Uncaptioned image] Vision: Diagram (2)
  • Text-only Version directly discards the diagram input from the Text-dominant version. Comparing this to the Text-lite version helps identify where MLLMs mainly obtain the contextual visual information for problem-solving, the Descriptive Information or diagram.
[…]
  • Vision-intensive Version further removes the Implicit Property from the Text-lite version. Without the strong visual condition in texts, MLLMs are challenged to intensively leverage their visual interpretation skills to gather sufficient cues for mathematical reasoning. The outcome demonstrates their proficiency in understanding mathematical relationships visually.
                        [Uncaptioned image] Text: EC + Question [Uncaptioned image] Vision: Diagram (4)
  • Vision-dominant Version, building upon the Text-lite version, excludes the Essential Condition from texts, instead annotating these measurements visually in diagrams. The textual content is narrowed down to Implicit Property and question statements. It demands MLLMs to recognize the Essential Condition exclusively from diagrams, and accurately correlate it with corresponding visual elements for problem-solving.
                        [Uncaptioned image] Text: IP + Question [Uncaptioned image] Vision: Diagram + EC (5)
  • Vision-only Version strips away the entire textual input, conveying the problem solely through the diagram. We render the remaining textual content in the Vision-dominant version onto the diagram, and minimize the text to an almost negligible prompt. This compels MLLMs to extract figures and recognize questions both from the visual information, standing as an ultimate test for their visual reasoning capabilities in mathematics.
[…]
Figure 6: Examples of the CoT Evaluation Strategy for MathVerse. We present two outputs from Qwen-VL-Max bai2023qwen with our CoT evaluation strategy, which assesses the fine-grained reasoning capabilities with a detailed explanation for error analysis.
2.3 CoT Evaluation Strategy
Compared to visual question-answering in general scenarios, the solving process of MLLMs for mathematical problems requires nuanced, step-by-step CoT reasoning. Considering two cases in Figure 6, one arrives at the correct solution albeit through incorrect intermediary steps, while the other demonstrates the opposite phenomenon. Therefore, the binary ‘Correct’ or ‘Incorrect’ evaluative approach of existing benchmarks is inadequate to examine the depth and precision of the multi-step reasoning process. To this end, we propose a CoT evaluation strategy to thoroughly assess their mathematical CoT skills in visual contexts, involving two prompting phases with GPT-4(V) openai2023gpt4v ; OpenAI2023GPT4TR .
Key-step Extraction.
Given the output of an MLLM, we first employ GPT-4, the language-only version, to extract $N$ pivotal steps within the reasoning sequence, denoted as $[s_{1},s_{2},\ldots,s_{N}]$, including the final answer $s_{A}$. Such key steps include significant computational outcomes, the identification of visual components, and critical immediate inferences. Note that, we only prompt GPT-4 with the MLLM’s output, deliberately omitting the original questions, diagrams, and ground-truth answers. This approach aims to mitigate the inherent bias of GPT-4 itself towards problem-solving and visual diagram interpretation, thereby concentrating solely on the logical coherence of the model output. In addition, we do not pre-define a ground-truth key-step template for each problem, but perform the extraction adaptively for the unique output of every MLLM. Since the problem potentially encompasses diverse possible solution pathways, and different MLLMs exhibit varying reasoning lengths and styles, the rigid template would harm the CoT evaluation accuracy.
Multi-step Scoring.
After the extraction phase, we utilize GPT-4V, the multi-modal version, to evaluate each critical step and culminate a comprehensive score. We feed the extracted key steps, the original questions, diagrams, and ground-truth answers all into GPT-4V, contributing to a holistic assessment, e.g., numerical computations, logical deductions, and visual interpretations. Therein, we observe that GPT-4V occasionally struggles with accurately recognizing elements within functional diagrams, leading to unstable evaluation for related problems. We thereby annotate additional information for function problems and together feed into GPT-4V, ensuring the quality of visual evaluation. Specifically, GPT-4V assesses each $N$ intermediate step with a binary score of ‘1’ (correct) or ‘0’ (incorrect), and derives the overall score by aggregating the correctness of the final answer. We formulate the scoring process as
$\text{Score}_{\text{final}}=\alpha\Big{(}\frac{1}{N}\sum_{i=1}^{N}\text{Score}(s_{i})\Big{)}+(1-\alpha)\text{Score}(s_{A}),\vspace{-0.2cm}$ (7)
where $\alpha$ denotes a balancing factor between the intermediate steps and the final answer $s_{A}$. We set $\alpha$ as 0.7 by default to underscore the significance of CoT reasoning. As exemplified in Figure 6, besides the fine-grained scoring, the CoT evaluation can also provide a detailed error analysis of each step, which is valuable and instructive for the development of MLLMs in the field.
3 Experiments
In this section, we conduct a systematic evaluation of existing Multi-modal Large Language Models (MLLMs) on MathVerse. We first introduce the experimental setup in Section 3.1. Then, we detail the quantitative results in Section 3.2 and narrate the error analysis in Section 3.3.
Table 2: Mathematical Evaluation on Six Problem Versions in MathVerse’s testmini Set. We calculate the ‘All’ score without averaging the ‘Text Only’ version. ‘CoT-E’ or ‘w/o’ denotes whether to employ the proposed CoT evaluation strategy or not. The highest accuracy for closed-source and open-source MLLMs is marked in red and blue respectively.
Model All Text Dominant Text Lite Text Only  Vision Intensive  Vision Dominant  Vision Only CoT-E w/o CoT-E w/o CoT-E w/o CoT-E w/o CoT-E w/o CoT-E w/o CoT-E w/o Baselines Random Chance - 12.4 - 12.4 - 12.4 - 12.4 - 12.4 - 12.4 - 12.4 Human - 64.9 - 71.2 - 70.9 - 41.7 - 61.4 - 68.3 - 66.7 LLMs ChatGPT ouyang2022training - - 51.3 33.3 38.5 18.9 51.3 33.3 - - - - - - GPT-4 OpenAI2023GPT4TR - - 63.4 46.5 40.7 20.7 63.4 46.5 - - - - - - Closed-source MLLMs Qwen-VL-Plus bai2023qwen 21.3 11.8 26.0 15.7 21.2 11.1 25.2 14.5 18.5 9.0 19.1 13.0 21.8 10.0 Gemini-Pro team2023gemini 35.3 22.3 39.8 27.6 34.7 23.7 44.5 27.9 32.0 19.4 36.8 20.3 33.3 20.5 Qwen-VL-Max bai2023qwen 37.2 24.8 42.8 30.3 37.7 24.8 47.9 32.2 33.6 20.6 35.9 23.3 35.9 25.1 GPT-4V openai2023gpt4v 54.4 38.3 63.1 52.1 56.6 40.9 60.3 46.1 51.4 34.9 50.8 33.6 50.3 29.8 Open-source MLLMs LLaMA-Adapter V2 gao2023llamaadapterv2 5.8 5.7 7.8 6.2 6.3 5.9 3.9 2.7 6.2 6.1 4.5 4.2 4.4 6.1 ImageBind-LLM han2023imagebind 10.0 9.3 13.2 11.4 11.6 11.3 12.9 11.7 9.8 8.9 11.8 11.2 3.5 3.4 mPLUG-Owl2 ye2023mplugowl2 10.3 4.6 11.6 6.6 11.4 6.3 13.8 6.1 11.1 6.3 9.4 5.6 8.0 4.9 MiniGPT-v2 chen2023minigpt 10.9 11.0 13.2 12.1 12.7 12.0 15.3 11.7 11.1 13.1 11.3 10.3 6.4 7.4 LLaVA-1.5 liu2023improvedllava 12.7 7.6 17.1 8.8 12.0 7.6 22.6 11.5 12.6 7.4 12.7 7.4 9.0 6.9 SPHINX-Plus gao2024sphinx 14.0 12.2 16.3 13.9 12.8 11.6 15.8 14.9 12.9 11.6 14.7 13.5 13.2 10.4 G-LLaVA gao2023g 15.7 16.6 22.2 20.9 20.4 20.7 21.6 21.1 16.5 17.2 12.7 14.6 6.6 9.4 LLaVA-NeXT liu2024llavanext 17.2 10.3 21.6 12.8 19.7 12.0 25.1 9.9 17.6 10.7 14.9 9.7 12.1 6.3 ShareGPT4V Chen2023ShareGPT4VIL 17.4 13.1 21.8 16.2 20.6 16.2 14.6 6.6 18.6 15.5 16.2 13.8 9.7 3.7 SPHINX-MoE gao2024sphinx 22.8 16.8 33.3 26.2 21.9 17.4 40.7 26.7 21.1 16.7 19.6 12.5 18.3 11.1 InternLM-XC2. dong2024internlm 25.9 16.3 36.9 20.2 28.3 14.3 42.5 24.5 20.1 14.2 24.4 17.5 19.8 15.2
3.1 Experimental Setup
Division of the testmini Subset.
MathVerse encompasses a comprehensive collection of 2,612 visual math problems, alongside 15,672 corresponding test instances. To enable faster evaluation and model development validation, we extract a smaller subset termed testmini including 788 problems and 4,728 instances. In constructing testmini, we employ a random sampling strategy across different subfields, maintaining a sample size proportional to the overall dataset to preserve its statistical representativeness. The remaining test set features 1,824 problems and 10,944 samples will be utilized for standard evaluation and publicly released in the future. In the subsequent experiments, all quantitative results are assessed using the testmini subset of MathVerse.
Evaluation Models.
We examine the performance of foundation models across three distinct categories on MathVerse: (a) Large Language Models (LLMs) as the text-only baseline, which only take textual questions as input, including ChatGPT OpenAI2023ChatGPT and GPT-4 OpenAI2023GPT4TR , (b) Closed-source MLLMs, represented by models like GPT-4V openai2023gpt4v , Gemini-Pro team2023gemini , Qwen-VL-Max bai2023qwen , and Qwen-VL-Plus, and (c) Open-source MLLMs, featuring models such as LLaVA-1.5 liu2023improvedllava (Vicuna-13B vicuna2023 ), LLaVA-NeXT liu2024llavanext (Vicuna-13B), SPHINX-MoE gao2024sphinx (Mixtral-8×7B albert24mixtral ), SPHINX-Plus (LLaMA2-13B touvron2023llama2 ), InternLM-XComposer2 dong2024internlm (InternLM2-7B team2023internlm ), LLaMA-Adapter V2 gao2023llamaadapterv2 (LLaMA-7B touvron2023llama ), ImageBind-LLM han2023imagebind (LLaMA-7B), MiniGPT-v2 chen2023minigpt (LLaMA2-7B), mPLUG-Owl2 ye2023mplugowl2 (LLaMA-7B), G-LLaVA gao2023g (LLaMA2-7B), and ShareGPT-4V Chen2023ShareGPT4VIL (Vicuna-13B).
Table 3: Mathematical Evaluation on Different Subjects and Subfields in MathVerse’s testmini Set. We report the scores averaging five problem versions except for the ‘Text Only’ version, and employ the CoT evaluation strategy by default. Len: Length; Anal: Analytic; Apply: Applied; Vol: Volume; Coord: Coordinate; Prop: Property; Exp: Expression; Apply: Applied. The highest accuracy for closed-source and open-source MLLMs is marked in red and blue respectively.
Model All Plane Geometry Solid Geometry Functions All Len Area Angle Anal Apply All Len Area Vol All Coord Prop Exp Apply Closed-source MLLMs Qwen-VL-Plus bai2023qwen 21.3 17.3 19.1 16.4 16.1 23.6 13.2 24.8 18.1 18.7 33.4 31.3 52.5 25.1 10.8 50.3 Gemini-Pro team2023gemini 35.3 33.0 32.2 42.6 28.4 30.2 32.3 33.4 35.0 29.3 36.1 28.3 25.7 26.6 10.8 51.3 Qwen-VL-Max bai2023qwen 37.2 38.4 41.7 46.4 32.6 40.6 38.7 33.7 25.4 28.3 42.6 38.4 43.7 35.5 13.6 61.0 GPT-4V openai2023gpt4v 54.4 56.9 60.8 63.4 52.6 48.5 60.9 50.2 54.8 39.9 56.8 52.8 72.3 47.1 30.9 70.1 Open-source MLLMs LLaMA-Adapter V2  gao2023llamaadapterv2 5.8 5.9 4.0 5.9 6.6 13.4 3.3 4.6 5.3 3.1 5.7 6.2 6.7 6.1 4.5 7.9 ImageBind-LLM han2023imagebind 10.0 9.7 12.1 9.9 9.2 10.2 4.8 4.6 4.9 3.5 5.3 14.9 12.3 13.8 4.6 25.9 mPLUG-Owl2 ye2023mplugowl2 10.3 7.7 8.2 6.0 5.7 12.4 10.6 11.0 9.2 6.7 15.7 17.4 22.8 18.6 5.3 22.2 MiniGPT-v2 chen2023minigpt 10.9 11.6 10.0 9.8 14.3 9.1 11.8 1.7 2.2 1.6 0.5 11.2 4.2 15.7 4.0 21.1 LLaVA-1.5 liu2023improvedllava 12.7 11.8 13.1 15.1 9.7 9.4 13.2 10.6 12.1 8.7 11.6 14.8 18.8 12.7 9.5 23.7 SPHINX-Plus gao2024sphinx 14.0 14.4 14.2 10.5 14.1 16.5 16.8 7.0 7.2 6.1 7.6 17.9 11.1 19.1 6.3 27.7 G-LLaVA gao2023g 15.7 20.2 17.3 13.6 26.5 5.9 23.1 5.0 10.3 4.4 3.1 9.2 9.1 9.1 1.3 15.5 LLaVA-NeXT liu2024llavanext 17.2 15.9 14.8 13.1 16.3 17.7 17.8 19.6 33.3 11.7 12.6 23.1 24.5 23.4 8.0 33.1 ShareGPT4V Chen2023ShareGPT4VIL 17.4 16.9 16.2 17.9 16.9 12.2 21.1 15.0 13.6 10.9 19.7 20.2 19.9 22.2 8.4 25.8 SPHINX-MoE gao2024sphinx 22.8 24.5 26.3 28.4 21.1 26.6 24.4 15.8 9.4 10.7 26.3 19.5 23.5 19.3 9.2 30.3 InternLM-XC2. dong2024internlm 25.9 26.2 27.1 29.7 20.6 18.5 22.2 20.1 34.5 14.1 25.2 23.7 24.4 24.9 10.6 36.3
Implementation Details.
All our experiments are conducted under a zero-shot setting, showcasing the generalization capacity of MLLMs for mathematical reasoning, without few-shot prompting or further fine-tuning. By default, we employ the Chain-of-Thought (CoT) prompting technique wei2022chain , which encourages MLLMs to perform complete reasoning steps for a fine-grained evaluation. A baseline representing random chance is established for comparison, for which we select one option at random for multiple-choice questions and utilize empty for free-form questions. In addition, we recruit ten qualified college students, and ask them to solve the problems in MathVerse independently, serving as a baseline for human performance. We conduct all experiments on NVIDIA A100 GPUs. As the text-only LLMs can only take text questions as input, we evaluate them with the first three problem versions, i.e., Text Dominant, Text Lite, and Text Only. For the ‘w/o’ results, we utilize the template in MathVista Lu2023MathVistaEM to prompt GPT-4 OpenAI2023GPT4TR for answer extraction, and directly score the final answer without the intermediate reasoning process.
3.2 Experimental Analysis
To best investigate the visual mathematical reasoning capabilities, we report the evaluation results of different models on MathVerse for the six transformed problem versions in Table 2 and twelve detailed subjects in Table 3. We mainly analyze the performance by the proposed Chain-of-Though (CoT) evaluation, and derive the following observations.
MLLMs Rely More on DI than Seeing Diagrams.
Comparing the Text-dominant and Text-only versions, with the elimination of visual input, most MLLMs even obtain an unexpected performance improvement, e.g., +5.1% for Qwen-VL-Max and +5.6% for InternLM-XComposer2. This suggests that the unsatisfactory visual encoding for mathematical diagrams instead severely harms the original problem-solving capacity of MLLMs. As exemplified in Figure 7, from the error analysis of our CoT evaluation strategy, we observe that Gemini-Pro can deduce the correct answer exclusively by the visual information within the Descriptive Information. Instead, the inaccurate visual perception of mathematical elements directly interferes with the outcome of problem-solving, turning correct answers into incorrect ones. In contrast, GPT-4V and ShareGPT-4V achieve better results in Text Dominant than in Text Only, indicating their relatively better visual encoding, which would not degrade the performance. However, they still encounter a larger performance drop by removing the redundant Descriptive Information than the diagram input, e.g., GPT-4V and ShareGPT-4V. This pattern demonstrates that they tend to capture more visual information for mathematical reasoning from the text content, instead of seeing the diagram itself.
Figure 7: A Typical Visual Perception Error by our CoT Evaluation Strategy. The example is an output from Gemini-Pro team2023gemini , where the correct reasoning of the Text-only version is distracted by the visual perception error within the diagram.
MLLMs are Moderately Effective at Perceiving IP.
By discarding the Implicit Property in question texts, a negligible decline in accuracy is noted from the Text-lite to Vision-intensive versions for most MLLMs. This is because the Implicit Property mainly encompasses the spatial layouts and geometric relationships, which demand minimal mathematical domain knowledge for interpretation. This outcome underscores the favorable visual perception skills of MLLMs for non-mathematical elements, which is not the primary obstacle hindering MLLMs in solving visual math problems.
MLLMs are Challenged to interpret EC from Diagrams.
Incorporating the Essential Condition within diagrams challenges MLLMs to accurately identify and understand these conditions in vision modality for mathematical problem-solving. Evidence from the Vision-dominant results indicates a notable decline in the performance of most MLLMs compared to the Text-lite accuracy, such as -5.8% for GPT-4V and -3.9% for InterLM-XComposer2. This reveals their inaccurate identification of mathematical symbols and an insufficient grasp of domain-specific knowledge required to associate identified measurements with relevant concepts.
MLLMs struggle to Solve Problems Entirely by Diagrams.
The scenario of Vision-only problems aligns more closely with real-world applications, where capturing an image is often more convenient than transcribing the problem into text. However, by rendering the whole question within the diagram, the mathematical problem-solving capacity of MLLMs is further diminished. This experiment unveils the great challenge for MLLMs to simultaneously understand mathematical conditions, questions, and figures from the visual input alone.
Closed-source MLLMs are Better-performed.
From the performance in both tables, we observe a consistently better performance achieved by closed-source MLLMs than open-sourced ones. Despite the gap with humans, GPT-4V attains the leading position among MLLMs, showcasing superior mathematical capabilities over problem versions and subjects, especially the challenging subfields like ‘Coord’ and ‘Prop’ (the property and coordinate solving of function problems). InternLM-XComposer2 and SPHINX-MoE are the best-performing open-source MLLMs, while still lagging behind Gemini-Pro with a margin of 9.4% and 12.5% overall accuracy, respectively, suggesting large improvement space.
LLMs Achieve Competitive Results to MLLMs.
Utilizing solely question texts as input, two LLMs, i.e., GPT-4 and ChatGPT, attain superior accuracy to most MLLMs in Text Dominant and Lite versions. Even in the absence of redundant Descriptive Information within Text-lite problems, GPT-4 outperforms InternLM-XComposer2 and SPHINX-MoE by substantial margins of 12.4% and 18.8%, respectively. These findings not only indicate the strong mathematical reasoning skills of LLMs, but further emphasize the deficiencies in diagram interpretation of existing MLLMs. Importantly, the performance of GPT-4 is only exceeded by GPT-4V, which demonstrates that a satisfactory diagram perception capability can enhance problem-solving for visual mathematics.
GPT-4(V) Beats Human in the Text-only Version.
Without the visual content provided in diagrams, human solvers often face challenges in deducing the correct answers due to the lack of sufficient information, e.g., 41.7% ‘w/o’ scores in Text-only problems. In contrast, GPT-4V and GPT-4 achieve the ‘w/o’ scores of 41.1% and 46.5%, respectively, which surpass the human performance. This comparison highlights their advanced reasoning capabilities in handling extreme scenarios, exhibiting more robustness for mathematical problem-solving given missing visual conditions.
Mathematical Training Benefits the Performance.
In addition to foundational visual instruction-following datasets, both SPHINX-MoE and InternLM-XComposer2 extend their training regimes to include specialized mathematical problems that are either text-only or visual, such as MathQA amini2019mathqa , Geometry3K Lu2021InterGPSIG , and MathInstruct yue2023mammoth . This approach of math-specific tuning contributes to their leading performance in MathVerse. Furthermore, G-LLaVA fine-tunes LLaVA-1.5 by a large-scale visual geometric dataset containing 170K enriched problems. This targeted refinement can improve several fields (‘Len’, ‘Angle’, and ‘Apply’) within the plane geometry subject. However, since G-LLaVA’s fine-tuning data does not include problems of analytic geometry, solid geometry, and functions, it harms the related results of LLaVA-1.5 due to catastrophic forgetting, e.g., -3.5% in ‘Anal’, -5.6% in ‘Solid Geometry’, and -5.6% in ‘Functions’. This phenomenon underscores the critical role of developing extensive, high-quality visual math data for effectively training MLLMs.
Discrepancy Between ‘CoT-E’ and ‘w/o’ Scores.
As illustrated by Table 2, the ‘CoT-E’ scores for MLLMs, in most cases, are much higher than ‘w/o’ scores, e.g., +16.1% for GPT-4V and +9.6% for InternLM-XComposer2. This observation demonstrates that our proposed CoT evaluation strategy identifies numerous correct intermediate reasoning steps, despite the final incorrect answer, highlighting the effectiveness of fine-grained assessment. In Figure 8, we present the statistics of variance between ‘CoT-E’ and ‘w/o’ scores within different MLLMs. Although GPT-4V attains top-tier performance, it exhibits a pronounced gap of 16.1% concerning the evaluation of CoT reasoning quality, similar to the 12.4% gap of Qwen-VL-Max. Conversely, SPHINX-MoE showcases favorable precision among open-source MLLMs, while preserving a relatively lower variance of two evaluation methods, i.e., 6.0% compared to InternLM-XComposer’s 9.6%. This indicates its consistent step-by-step reasoning throughout the problem-solving process.
Figure 8: Results with and without CoT Evaluation in MathVerse. Referring to Table 2, we denote the ‘w/o’ results in blue pillars, and highlight the increase and decrease magnitude with ‘CoT-E’ by green and red colors, respectively.
3.3 Error Analysis
Figure 9: Distribution of GPT-4V’s openai2023gpt4v Errors in Reasoning and Answers. For the six problem versions in MathVerse, we provide the statistics of errors made by GPT-4V based on their occurrence in answers (‘Ans.’) and reasoning processes (‘Rea.’).
To delve into the fine-grained predictions, we select the best-performing MLLM, GPT-4V openai2023gpt4v , to understand its modes of success and failure. Our proposed CoT evaluation strategy has produced a detailed assessment of model output, including step-wise scores and explanation, reducing extensive manual effort in identifying and analyzing errors. We conduct our analysis on the two-step output from the CoT evaluation across the entire dataset, focusing on two key dimensions.
Errors in Reasoning or Answer?
In Figure 9, we showcase the statistics of different error distributions in six problem versions of MathVerse. We define the following six error categories: correct final answer with correct/partially correct/incorrect CoT reasoning and incorrect final answer with correct/partially correct/incorrect CoT reasoning. For all six versions, the incorrect final answers are mostly caused by the partially incorrect reasoning process. In addition, a number of problems with correct answers are accompanied by partially or entirely incorrect reasoning, e.g., 15.3% in Text Dominant, which cannot be detected by the traditional True or False evaluation. As we remove the content within textual questions and enrich the visual diagram, e.g., from Text Dominant and Lite to Vision Dominant and Only, we observe a progressive increase in the error rate of ‘incorrect final answer with incorrect CoT reasoning’, indicating that MLLMs are challenged to conduct high-quality intermediate reasoning by capturing more information from the visual input.
Figure 10: Distribution of GPT-4V’s openai2023gpt4v Errors within Different Types. We present the statistics of four error types by GPT-4V in the six problem versions, i.e., Visual Perception Error, Reasoning Error, Calculation Error, and Knowledge Error.
What Types of Errors?
To further investigate the specific error types, we survey the problems with errors that occur either within the reasoning process or the final answer. As depicted in Figure 10, we divide the errors of GPT-4V into four distinct types: visual perception error, reasoning error, knowledge error, and calculation error. Consistent with our findings in the main paper, the primary source of errors in problem-solving attributes to the inaccurate interpretation of mathematical diagrams, which significantly impedes the performance of MLLMs. For the problem versions that demand advanced diagram interpretation, e.g., Vision Dominant and Only, we observe a notable increase in the rate of visual perception errors, demonstrating an urgent need for stronger visual encoders in MLLMs. Moreover, reasoning errors also account for a considerable percentage, indicating that the logical deduction skills of MLLMs still require improvement. As expected, knowledge errors do not significantly hinder the mathematical reasoning capabilities of MLLMs in MathVerse.
4 Conclusion
In this paper, we propose a comprehensive and specialized benchmark, MathVerse, for the visual mathematical problem-solving capacity of MLLMs. We meticulously collect high-quality math problems with diagrams spanning three primary subjects and twelve subfields. Given the issues within current benchmarks, we transform each problem into six versions, investigating whether and how much MLLMs can interpret the visual math diagrams. We also propose a CoT evaluation strategy for finer-grained assessment of the intermediate reasoning process of MLLMs. By evaluating various closed-source and open-source models, MathVerse unveils that most existing MLLMs struggle to accurately understand mathematical diagrams, and even attain higher results without visual input. This indicates the potential of developing more advanced math-specific vision encoders for stronger multi-modal mathematical reasoning.
Appendix Overview
  • Section A: Related work.
  • Section B: Additional experimental details.
  • Section C: More dataset details.
  • Section D: Comparison to current benchmarks.
  • Section E: Limitation and future work.
  • Section F: Qualitative examples.
Appendix A Related Work
Multi-modal Large Language Models (MLLMs),
building upon the prevalence of Large Language Models (LLMs) touvron2023llama ; touvron2023llama2 ; OpenAI2023ChatGPT ; albert24mixtral ; brown2020language and large vision models Radford2021LearningTV ; kirillov2023segment ; zhang2023personalize ; zhang2023prompt ; zhang2023learning , have become increasingly prominent in the field. They extend LLMs to tackle a diverse range of tasks and domains, including the mainstream 2D images li2022blip ; instructblip ; alayrac2022flamingo ; li2023mimic and other modalities, such as 3D point clouds guo2023point ; xu2023pointllm ; hong20243d , audio han2023imagebind ; su2023pandagpt , and video zhang2023video ; chen2023videollm . Noteworthy examples like OpenAI’s GPT-4V openai2023gpt4v and Google’s Gemini team2023gemini exhibit exceptional visual understanding and reasoning capabilities, setting new benchmarks in multi-modal performance. However, their closed-source nature poses a barrier to the broader application and development of MLLMs. Concurrently, another line of work is dedicated to exploring advanced MLLMs open-source to the community. Prior efforts like LLaMA-Adapter zhang2024llamaadapter ; gao2023llamaadapterv2 , LLaVA liu2023llava ; liu2024llavanext ; liu2023improvedllava , and MiniGPT-4 zhu2023minigpt ; chen2023minigpt leverage a frozen CLIP Radford2021LearningTV model for image encoding, and inject the visual cues into LLaMA touvron2023llama for multi-modal instruction tuning. The subsequent mPLUG-Owl ye2023mplugowl ; ye2023mplugowl2 , Qwen-VL bai2023qwen , InternLM-XComposer dong2024internlm , and SPHINX lin2023sphinx ; gao2024sphinx further push the frontier of MLLMs in understanding and generalizing across visual contexts. Despite comprehensive benchmarks fu2023mme ; liu2023mmbench ; Li2023SEEDBenchBM ; xu2023lvlm on general visual instruction-following scenarios, the specific potential of MLLMs for visual mathematical problem-solving remains under-explored. In this paper, we introduce the MathVerse benchmark to comprehensively evaluate the visual mathematical reasoning and diagram understanding skills of MLLMs, providing unique perspectives for future research directions.
Mathematical Reasoning Benchmarks
have emerged as a significant area of focus, posing considerable challenges for large foundational models, e.g., LLMs and MLLMs. Initially, datasets in this realm are designed to address basic algebraic hendrycksmath2021 and arithmetic Roy2016SolvingGA word problems, which are relatively limited in scope and volume. Subsequent efforts, including MATH hendrycksmath2021 , GSM8K cobbe2021training , and MMLU hendryckstest2021 , expand the range and quality of textual mathematical problems. These datasets feature a broader spectrum of difficulties, establishing a robust benchmark for the evaluation of general and math-specific LLMs zhou2023solving ; yue2023mammoth ; wang2024mathcoder ; gao2023g ; luo2023wizardmath . Besides the text-only assessment, there is a growing demand for comparable, high-quality benchmarks for evaluating mathematical problem-solving in visual contexts, with the rapid progress of MLLMs. There are prior attempts, such as GeoQA Chen2021GeoQAAG , UniGeo Chen2022UniGeoUG , and Geometry3K Lu2021InterGPSIG , which focused exclusively on geometric problems. The recently proposed MathVista Lu2023MathVistaEM broadens the scope to incorporate a variety of multi-modal tasks involving mathematical reasoning, and MMMU yue2023mmmu covers college-level questions demanding intricate, domain-specific knowledge. However, our analysis identifies three main shortcomings within the current visual math benchmarks, as elaborated in Section 1 of the main paper. Therefore, we propose MathVerse specialized in the multi-modal mathematical evaluation of MLLMs, comprising twelve subjects, six problem versions, and 20K test samples. Our objective is to thoroughly investigate whether and how much MLLMs genuinely interpret visual diagrams for mathematical reasoning.
Table 4: The Release Time and Model Source of MLLMs Used in MathVerse.
Model Release Time
Source
ChatGPT ouyang2022training 2022-11
https://platform.openai.com/
GPT-4 OpenAI2023GPT4TR 2023-03
https://platform.openai.com/
Qwen-VL-Plus bai2023qwen 2023-11
https://help.aliyun.com/zh/dashscope/developer-reference/vl-plus-quick-start
Gemini-Pro team2023gemini 2023-12
https://ai.google.dev/
Qwen-VL-Max bai2023qwen 2024-01
https://help.aliyun.com/zh/dashscope/developer-reference/vl-plus-quick-start
GPT-4V openai2023gpt4v 2023-09
https://platform.openai.com/
LLaMA-Adapter V2 gao2023llamaadapterv2 2023-04
https://github.com/OpenGVLab/LLaMA-Adapter/tree/main/llama_adapter_v2_multimodal7b
LLaVA-1.5 liu2023improvedllava 2023-10
https://huggingface.co/liuhaotian/llava-v1.5-13b
MiniGPT-v2 chen2023minigpt 2023-10
https://github.com/Vision-CAIR/MiniGPT-4
mPLUG-Owl2 ye2023mplugowl2 2023-11
https://huggingface.co/MAGAer13/mplug-owl2-llama2-7b
G-LLaVA gao2023g 2023-12
https://github.com/pipilurj/G-LLaVA/tree/main
ImageBind-LLM han2023imagebind 2023-05
https://github.com/OpenGVLab/LLaMA-Adapter/tree/main/imagebind_LLM
ShareGPT4V Chen2023ShareGPT4VIL 2023-11
https://huggingface.co/Lin-Chen/ShareGPT4V-13B
SPHINX-Plus lin2023sphinx 2023-11
https://huggingface.co/Alpha-VLLM/LLaMA2-Accessory/tree/main/finetune/mm/SPHINX/SPHINX-v2-1k
LLaVA-NeXT liu2024llavanext 2024-01
https://huggingface.co/liuhaotian/llava-v1.6-vicuna-13b
SPHINX-MoE gao2024sphinx 2024-01
https://huggingface.co/Alpha-VLLM/LLaMA2-Accessory/tree/main/finetune/mm/SPHINX/SPHINX-MoE
InternLM-XComposer2 dong2024internlm 2024-01
https://huggingface.co/internlm/internlm-xcomposer2-vl-7b
Appendix B Additional Experimental Details
Model Sources.
For different MLLMs, we select their latest models and best-performing configurations for evaluation to fully reveal their visual mathematical proficiency. Table 4 presents the release time and model sources of MLLMs used in MathVerse.
Table 5: Input Prompt of MLLMs for Response Generation. We adopt two different prompts for the free-form and multiple-choice questions. Note that these prompts are used for five problem versions except for the Vision-only version.
Question Prompt
Free-form Question
Please first conduct reasoning, and then answer the question and provide the final value, e.g., 1, 2.5, 300, at the end.
Question: $\{question\}$
Multiple-choice Question
Please first conduct reasoning, and then answer the question and provide the correct option letter, e.g., A, B, C, D, at the end.
Question: $\{question\}$
Table 6: Input Prompt for Vision-only Problems. Especially for the Vision-only version without textual input, we add “According to the question shown in the image” at the beginning of the prompt, and remove the “Question:” at the end.
Question Prompt
Free-form Question
According to the question shown in the image, please first conduct reasoning, and then answer the question and provide the final value, e.g., 1, 2.5, 300, at the end.
Multiple-choice Question
According to the question shown in the image, please first conduct reasoning, and then answer the question and provide the correct option letter, e.g., A, B, C, D, at the end.
Prompt for Response Generation.
We adopt two types of prompts respectively for the free-form and multiple-choice questions, as shown in Table 5. We inspire the Chain-of-Thought (CoT) reasoning capabilities of MLLMs by using the phrase “first conduct reasoning”. Especially for the Vision-only problem version in Table 6, we add “According to the question shown in the image” at the beginning to remind MLLMs to read the questions rendered within diagrams, where the textual input for MLLMs only contains the prompt itself.
Figure 11: Manual Annotations for Function Problems in MathVerse. We provide detailed annotations, e.g., function expression and properties, for the diagrams of 534 function problems, which benefits the accuracy of GPT-4V openai2023gpt4v for CoT evaluation.
Prompt for the CoT Evaluation.
Our proposed CoT evaluation contains two steps, i.e., key-step extraction and multi-step scoring, which prompt GPT-4 OpenAI2023GPT4TR and GPT-4V openai2023gpt4v , respectively. The input configuration is listed in Table 7. We utilize the text-only GPT-4 in the first step to extract multiple key steps within the model’s unstructured output, without feeding the question information. In the second step, we input the extracted key-step reasoning and all the available content related to the problem into GPT-4V, allowing for a holistic assessment, including diagram interpretation, logical reasoning, and numerical computation. In Figure 11, we showcase the manual annotation for critical information within functional diagrams, e.g., function expression and properties. This assists GPT-4V in evaluating the visual perception accuracy of MLLMs for function graphs.
Table 7: Configuration for the CoT Evaluation Strategy. We conduct two evaluation phases, respectively by prompting the text-only GPT-4 OpenAI2023GPT4TR and GPT-4V openai2023gpt4v . The symbol ‘XXX’ denotes the given one-shot sample, abbreviated for brevity. The ‘Annotation’ (represented in grey) in the second phase is only required for function problems.
Phase Input Prompt
Key-step Extraction (GPT-4) Model Output
I will give you a detailed solving procedure or a single answer for a math problem.
If it is a procedure, you need to extract the key solution steps and list them accordingly in markdown syntax. If it is just a single answer, output the answer directly.
Here are examples:
– Model output: XXX
– Extracted: 1. XXX 2. XXX 3. XXX
– Model output: 2.2
– Extracted: The single answer is 2.2
Here is what you need to extract:
Model output: $\{model\ output\}$
Extracted:
Multi-step Scoring (GPT-4V) Extracted Steps Question Diagram Answer Annotation
I will first give you a visual math problem, including the question, diagram, ground-truth answer, and detailed annotation of the diagram, and then give you a model output containing multiple key solution steps.
Please think step by step and output the Average score, along with the Final answer score in the end, as described below:
– Average score: Evaluate, based on the given question, answer, diagram, and diagram annotation, whether each solution step is correct in logical reasoning, visual perception, and numerical computation, with an incorrect score of 0 and a correct score of 1. Then, calculate the average score of multiple steps.
– Final answer score: Match the model’s final answer with the ground truth answer, scoring 1 if it matches and 0 if it doesn’t.
– If the model output only includes a single step or answer, the Average score and Final answer score are the same.
Question: $\{question\}$
Ground-truth answer: $\{answer\}$
Diagram annotation: $\{annotation\}$
Model output: $\{extracted\ steps\}$
Average score:
Final answer score:
Human Performance Assessment.
We recruit ten qualified college students specifically for the evaluation of human performance on MathVerse. These individuals are kept separate from the data curation stage, eliminating the possibility of them encountering the solutions beforehand. We allocate to each student the questions from a specific problem version. This strategy is to prevent them from gaining additional information from another version to answer questions, e.g., leveraging the textual Implicit Property from the Text-lite version to solve Text-intensive problems. They are asked to directly provide the final answer without detailed reasoning. Therefore, we do not report the CoT evaluation results for human performance, alongside the ‘Random Chance’ baseline.
[…]
Appendix C More Dataset Details
C.1 Data Curation
This paper engages twelve expert annotators for data curation, consisting of senior undergraduate and graduate students from across the globe with a strong background in science. In collaboration with the authors, they are required to mainly complete five tasks concerning data collection, categorization, quality review, problem version transformation, and function diagram annotation.
Data Collection.
[…]
Data Categorization and Review.
We first ask the human annotators to categorize the problems into three primary subjects, i.e., plane geometry, solid geometry, and functions. Within each subject, according to the definitions in Section C.2, the math problems are further divided into twelve fine-grained categories. At the same time, we meticulously review the collected dataset. We manually rectify the problems with incorrect answers and discard the problems with multiple diagrams, visual solutions, and too much similar content to others. Finally, 2,612 high-quality math problems with paired diagrams are preserved for MathVerse, spanning diverse subjects and subfields.
Figure 12: Manual Annotations for Descriptive Information in MathVerse. For some collected problems, we are required to supplement additional Descriptive Information (highlighted in red) to distinguish the Text-dominant version.
Figure 13: Manual Modification for Textual Essential Condition in MathVerse. For the original problems shown, we transfer some of the Essential Condition from diagrams to question texts (highlighted in green) to mark the Vision-dominant version.
Transformation of Problem Versions.
Given the three types of textual information within questions, human annotators rigorously transform each problem into six different versions as discussed in Section 2.2 of the main paper. We utilize Microsoft PowerPoint to annotate the diagrams in the Vision-dominant version, and employ Matplotlib to render the questions onto the diagrams in the Vision-only version. As illustrated in Figure 12, for problems with minimal Descriptive Information, we manually enhance the question text with additional contextual description about the diagram to differentiate the Text-dominant version. In the case of questions in Figure 13, where the Essential Condition has been fully depicted in the diagrams, we remove some of this content from the diagram and incorporate it into the text to mark the Vision-dominant version.
C.2 Subject and Subfield Definition
The visual math problems within MathVerse encompass three primary subjects, plane geometry, solid geometry, and functions, alongside twelve finer-grained subfields, which comprehensively evaluate the diagram understanding and mathematical reasoning capabilities of MLLMs.
Plane Geometry
is a fundamental area that explores the properties and relations of points, lines, and surfaces in a two-dimensional plane. This subject delves into concepts such as angles, triangles, circles, and polygons, offering a rich context for assessing the spatial comprehension and logical deduction skills of MLLMs. We divide it into five subfields, as exemplified in Figure 14:
  • Length focuses on the measurement and comparison of distances between points. This subfield includes understanding the properties of lines, segments, and their use in determining the perimeters of geometric shapes, which is foundational for MLLMs to solve plane geometry problems.
  • Area examines the size of two-dimensional surfaces. It encompasses calculating the areas of various shapes, such as triangles, rectangles, circles, and more complex polygons, by applying specific formulas and principles, which is crucial for comprehending the concept of space within geometry.
  • Angle involves the study of angles and their properties, including different types of angles (acute, right, and obtuse), angle measurement, and the relationships between angles, particularly in polygons. This subfield demands the advanced spatial perception capacity of MLLMs.
  • Analytic Geometry, also known as coordinate geometry, merges algebra and geometry to solve geometric problems using coordinate systems, exploring the calculation and reasoning of equations for geometric shapes. MLLMs are evaluated on their coordinate identification and algebraic capabilities.
  • Applied Geometry relate to the application of geometric principles to solve real-world and theoretical problems. It challenges MLLMs to first understand the background information within questions, and apply their knowledge of lengths, areas, angles, and analytic geometry for problem-solving.
Figure 14: Examples of Five Subfields in Plane Geometry, spanning Length, Area, Angle, Analytic, and Applied Geometry problems. We showcase the Text-lite version.
Solid Geometry
focuses on the study of three-dimensional objects that have depth, length, and width, thereby offering a more complex and enriched exploration of spatial structures. This subject investigates a variety of shapes such as cubes, cylinders, spheres, and pyramids, and assesses MLLMs to tackle questions concerning the volume, surface area, and geometric properties of these solids. This subject contains three subfields, as exemplified in Figure 15:
  • Length, extending from the 2D counterpart, focuses on measuring the edges and curves that define three-dimensional objects. It involves determining the linear distance between points in space, the perimeters of bases of solids, and the height or depth of objects. This measurement is a foundational element for MLLMs in analyzing geometric solids.
  • Area encompasses the calculation of the total area covered by the outer surfaces of solids. This normally requires MLLMs to break down complex shapes into several simpler components for area calculation in plane geometry, assessing their spatial and logical reasoning performance.
  • Volume pertains to measuring the space enclosed within three-dimensional objects. This demands MLLMs to precisely identify the geometric solids and apply accurate formulas to calculate the volume, which evaluates their mathematical knowledge application and calculation skills.
Functions
involve analyzing mathematical functions to understand the relationship between variables. These challenges range from simple tasks, like calculating a function value for a given input, to more complex scenarios, such as exploring the behavior and representation of various function types. We evaluate MLLMs by four types of function problems, exemplified in Figure 16:
Figure 15: Examples of Three Subfields in Solid Geometry, spanning Length, Area, and Volume problems. We showcase the Text-lite version.
  • Function Coordinate focuses on interpreting and extracting coordinate-level information from graphical representations of functions. It includes tasks such as identifying specific coordinate values of points on the graph and observing intersection points between functions and axes, which test the MLLM’s basic proficiency in functional visual perception.
  • Function Property emphasizes the model’s capacity to discern and deduce the inherent properties of functions from their graphs, such as symmetry, asymptotes, extrema (maximum and minimum points), and intervals of increase or decrease. These problems can reveal the understanding of MLLMs for the deeper characteristics of functions.
  • Function Expression refers to the direct analysis using the algebraic expressions of functions, widely including linear, quadratic, polynomial, exponential, logarithmic, and piece-wise functions. It challenges MLLMs to extract specific function expressions and apply transformations, bridging the gap between abstract mathematical reasoning and visual interpretation.
  • Applied Function, similar to the applied geometry problems, requires MLLMs to leverage their functional knowledge and theorems in real-world scenarios, e.g., modeling economic data, predicting physical phenomena, and calculating probabilities. This assesses the MLLM’s capabilities to understand functions in both theoretical and practical contexts.
Figure 16: Examples of Four Subfields in Functions, spanning Function Coordinate, Property, Expression, and Applied problems. We showcase the Text-lite version.
C.3 Detailed Statistics of MathVerse
More Data Statistics.
In Table C.3, we provide a more detailed data statistics of MathVerse. Therein, the 534 newly annotated questions refer to all the function problems, for which we meticulously annotate critical functional information, as depicted in Figure 11. The number of newly annotated diagrams represents the 5,224 math problems in the Vision-dominant and Vision-only versions. For these problems, we respectively integrate the Essential Condition and all textual content with the diagrams. We also list the numbers of multiple-choice answers, where A, B, C, and D are almost uniformly distributed.
Table 8: Statistics of MathVerse.
Table 9: Length of Different Problem Versions in MathVerse.
Statistic Number Total questions 2,612  - Subjects/subfields 3/12  - Multiple-choice questions 1,631 (62.4%)  - Free-form questions 981 (37.6%)  - Newly collected questions 1,236 (47.3%)  - Existing-dataset questions 1,376 (52.7%)  - Questions with explanations 1,236 (47.3%)  - Newly annotated questions 534 (20.4%) Multiple-choice question  - Proportion of answer A 585 (22.4%)  - Proportion of answer B 828 (31.7%)  - Proportion of answer C 703 (26.9%)  - Proportion of answer D 444 (17.0%)  - Proportion of answer E&F 52 (2.0%) Total test samples 15,672  - Newly annotated samples 10,448 (66.7%)  - Newly annotated diagrams 5,224 (33.3%)  - Samples of each version 2,612 (16.7%) Number of unique images 2,420 (92.6%) Number of unique questions 2,573 (98.5%) Number of unique answers 847 (32.4%)
Problem Version    Word    Character Text Dominant & Text Only  - Maximum question length 203 1,311  - Maximum answer length 17 102  - Average question length 35.7 204.8  - Average answer length 1.4 6.3 Text Lite  - Maximum question length 179 1,173  - Maximum answer length 17 102  - Average question length 22 133.8  - Average answer length 1.4 6.3 Vision Intensive  - Maximum question length 171 1,126  - Maximum answer length 17 102  - Average question length 18.8 116.8  - Average answer length 1.4 6.3 Vision Dominant  - Maximum question length 176 1,132  - Maximum answer length 17 102  - Average question length 17.6 123.5  - Average answer length 1.4 6.3
Table 9: Length of Different Problem Versions in MathVerse.
Problem Length Variance.
In Table C.3, we highlight the variance in question and answer lengths across the five problem versions in MathVerse, excluding the Vision-only category due to its absence of text. For both word and character levels, as we remove the pre-defined textual elements (Descriptive Information, Implicit Property, and Essential Condition), the maximum and average lengths of questions decrease accordingly, while the answer lengths remain the same. In Figure 17, we visualize the word-level variation of question length for the three problem versions: Text Dominant (blue), Text Lite (green), and Vision Dominant (red). By progressively omitting Descriptive Information and Essential Condition from the Text-dominant version, we observe a clear downward trajectory for the question length distribution and their average values.
Figure 17: Distribution of Question Length for Three Problem Versions. We exclude the Descriptive Information and Essential Condition from the Text-dominant problems, respectively creating the Text-lite and Vision-dominant versions.
Appendix D Comparison to Current Benchmarks
In this section, we offer a detailed comparison between MathVerse and existing multi-modal mathematical benchmarks, i.e., geometry-specific benchmarks chen2021geoqa ; cao2022augmented ; chen2022unigeo ; lu2021inter ; seo2015solving , MathVista Lu2023MathVistaEM , and MMMU yue2023mmmu , from the following four aspects:
The Investigation of Diagram Interpretation Capacity.
As discussed in Figure 1 of the main paper, the math problems in most existing datasets contain excessive redundant information in textual content, which is repetitive to the visual elements in diagrams. This issue enables MLLMs to potentially bypass the process of visual understanding, and thereby cannot determine whether and how much MLLMs truly interpret the math diagram. In contrast, our MathVerse includes six problem versions with different information content across text and vision. By comparing the performance variance between different problem versions, we can thoroughly investigate the mathematical diagram interpretation capabilities of MLLMs for the first time.
Figure 18: Diagram Examples of Math-related Tasks in MathVista Lu2023MathVistaEM . These tasks are not strongly correlated to the mathematical reasoning skills of MLLMs, probably skewing the assessment emphasis towards visual math problems.
Figure 19: Diagram Examples with Required Knowledge in MMMU yue2023mmmu . These math problems demand MLLMs to comprehend college-level domain knowledge, potentially hindering them from fully exerting mathematical reasoning skills.
Evaluation Approach.
Previous benchmarks adopt a simple True or False metric to score the response from MLLMs, which lacks fine-grained information and intermediate reasoning assessment, as analyzed in Figure 2 of the main paper. In contrast, MathVerse adopts a unique CoT evaluation strategy by examining each crucial solution step within the model output. This approach not only unveils the CoT reasoning quality of MLLMs, but also provides detailed error analysis, serving as valuable guidance for future enhancement.
The Depth and Width in Math Problems.
The geometry-specific benchmarks evaluate only a limited dimension of mathematical skills in MLLMs. MathVista instead incorporates a variety of math-related question-answering tasks, e.g., textbook figures, tables, plots, charts, puzzles, and synthetic scenes, as exemplified in Figure 18. However, the integration of these peripheral tasks (covering more than 70%) might divert the focus from the specialized mathematical evaluation of MLLMs. In addition, MMMU focuses on college-level complexity, requiring advanced domain-specific knowledge, as depicted in Figure 19. Given this, the lack of profound mathematical theorems would restrict the performance of MLLMs, biasing the evaluation of logical reasoning and visual perception proficiency. Therefore, our MathVerse concentrates on specialized visual math problems (plane geometry, solid geometry, and functions) with a moderate difficulty (high-school level), aiming to fully exert the capabilities of MLLMs.
Total Volume of Test Samples.
We summarize the size of test instances for different datasets in Table 10. As demonstrated, our MathVerse offers a considerably larger number of samples than others, nearly three times to MathVista and twenty times to GeoQA+, including meticulously annotated six versions of visual math problems. This contributes to a comprehensive and robust evaluation of visual mathematical reasoning capabilities.
Table 10: Number of Test Samples in Different Benchmarks.
Benchmark
GEOS
Geo3K
GeoQA+
MathVista
MMMU-Math
MathVerse
Test Samples
119
601
755
6,141
540
15,672
Appendix E Limitation and Future Work
While our MathVerse takes a step forward in the field of visual mathematical evaluation for MLLMs, it is important to recognize several limitations as follows.
We have categorized the math problems in MathVerse by various criteria, including subjects, subfields, and versions featuring differing degrees of multi-modal content. These categorization approaches evaluate the capabilities of MLLMs from multiple dimensions. Nevertheless, it is also meaningful to further divide the problems based on their difficulty levels, akin to MATH hendrycksmath2021 , a text-only benchmark defining five levels of difficulty. This additional layer of differentiation can provide deeper insights into the problem-solving abilities of MLLMs across a spectrum of challenges, which we leave as future work.
The curated dataset in MathVerse focuses on math problems in the high school level with moderate difficulty, which aims to fully demonstrate the mathematical reasoning skills within current MLLMs. However, with the advancement of architecture and training methodologies, future MLLMs have the potential to grasp more complex knowledge and theorems across a variety of domains. Therefore, there is significant value in further augmenting MathVerse with problems spanning broader complexity and disciplines, including those at the college level and within scientific fields. By transforming the expanded problems into different versions, we can facilitate a more comprehensive and robust evaluation of MLLMs for their diagram interpretation and reasoning capabilities.
Moreover, the problems in MathVerse and other current mathematical benchmarks are mainly in English. Given that some multilingual MLLMs han2023imagebind ; bai2023qwen have been developed, existing evaluation cannot reveal their full capabilities when confined to a single language. The incorporation of multilingual visual math problems would not only extend the dataset’s global applicability, but also enhance the assessment of MLLMs for linguistic diversity and understanding.
Appendix F Qualitative Examples
To ease the understanding, we offer a variety of qualitative examples in MathVerse. In Section F.1, we showcase the meticulously transformed six versions of visual math problems. In Section F.2, we compare the response of different MLLMs on Text-lite problems, including GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx . Specifically, we present the key-step extraction output by the CoT evaluation, and mark the multi-step scoring results aside. In Section F.3, we provide the response comparison of GPT-4V for three problem versions in MathVerse, i.e., Text Dominant, Text Lite, and Vision Dominant.
F.1 Comparison of Six Problem Versions
Please refer to Figures 13$\sim$15.
Figure 20: Comparison of Six Problem Versions in MathVerse.
Figure 21: Comparison of Six Problem Versions in MathVerse.
Figure 22: Comparison of Six Problem Versions in MathVerse.
F.2 Response of Different MLLMs
Please refer to Figures 16$\sim$21.
Figure 23: Response Comparison of GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx . We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
Figure 24: Response Comparison of GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx . We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
Figure 25: Response Comparison of GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx . We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
Figure 26: Response Comparison of GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx . We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
Figure 27: Response Comparison of GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx . We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
Figure 28: Response Comparison of GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx . We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
F.3 Response of Different Problem Versions
Please refer to Figures 22$\sim$27.
Figure 29: Response of Different Problem Versions by GPT-4V openai2023gpt4v . By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams. We only showcase the Text-dominant version of the problem for brevity.
Figure 30: Response of Different Problem Versions by GPT-4V openai2023gpt4v . By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams. We only showcase the Text-dominant version of the problem for brevity.
Figure 31: Response of Different Problem Versions by GPT-4V openai2023gpt4v . By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams. We only showcase the Text-dominant version of the problem for brevity.
Figure 32: Response of Different Problem Versions by GPT-4V openai2023gpt4v . By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams. We only showcase the Text-dominant version of the problem for brevity.
Figure 33: Response of Different Problem Versions by GPT-4V openai2023gpt4v . By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams. We only showcase the Text-dominant version of the problem for brevity.
Figure 34: Response of Different Problem Versions by GPT-4V openai2023gpt4v . By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams. We only showcase the Text-dominant version of the problem for brevity.

### v2

@@
 1 Introduction
-With the substantial advances of big data and computational power, Large Language Models (LLMs) brown2020language ; albert24mixtral ; touvron2023llama ; touvron2023llama2 ; vicuna2023 , such as ChatGPT OpenAI2023ChatGPT and GPT-4 OpenAI2023GPT4TR , have emerged as a central point of interest in both industry and academia.
-To broaden their applicability across diverse contexts, Multi-modal Large Language Models (MLLMs) zhang2024llamaadapter ; gao2023llamaadapterv2 ; su2023pandagpt ; chen2023minigpt ; ye2023mplugowl have recently become a fast-evolving track, exemplified by the latest GPT-4V openai2023gpt4v , Gemini team2023gemini , and the open-source LLaVA liu2023llava ; liu2023improvedllava ; liu2024llavanext and SPHINX lin2023sphinx ; gao2024sphinx .
-Concurrently, a diverse array of evaluation benchmarks fu2023mme ; liu2023mmbench ; Li2023SEEDBenchBM ; fu2023challenger ; sun2024journeydb are curated to assess their visual comprehension performance across different domains.
+With the substantial advances of big data and computational power, Large Language Models (LLMs) [4, 28, 55, 56, 13], such as ChatGPT [45] and GPT-4 [46], have emerged as a central point of interest in both industry and academia.
+To broaden their applicability across diverse contexts, Multi-modal Large Language Models (MLLMs) [66, 20, 52, 11, 61, 70] have recently become a fast-evolving track, exemplified by the latest GPT-4V [47], Gemini [22], and the open-source LLaVA [39, 34, 32, 30] and SPHINX [36, 21].
+Concurrently, a diverse array of evaluation benchmarks [17, 40, 33, 18, 53] are curated to assess their visual comprehension performance across different domains.
 Notably, the capability to solve mathematical problems involving diagrams serves as a critical measure, offering insights into the multi-modal logical thinking prowess of MLLMs.
 This task demands MLLMs to accurately decode the visual elements within input diagrams (characters and figures), and correlate them with the condition specified by textual questions for mathematical reasoning.
-Previous efforts lu2021inter ; seo2015solving , e.g., GeoQA chen2021geoqa ; cao2022augmented and UniGeo chen2022unigeo , concentrate on the challenging geometric problems, while the recent MathVista Lu2023MathVistaEM and MMMU yue2023mmmu expand the scope to encompass broader disciplines, including functions, charts, and scientific problems.
-Figure 1: (a) We showcase three examples of Text Redundancy (highlighted in red) within existing visual math benchmarks chen2021geoqa ; Lu2023MathVistaEM ; yue2023mmmu .
-(b) We report an ablation study by respectively removing the redundant texts and input diagrams on 120 randomly selected problems, for closed-sourced openai2023gpt4v ; team2023gemini ; bai2023qwen and open-sourced gao2024sphinx ; liu2024llavanext ; dong2024internlm MLLMs.
+Previous efforts [42, 51], e.g., GeoQA [9, 5] and UniGeo [7], concentrate on the challenging geometric problems, while the recent MathVista [41] and MMMU [63] expand the scope to encompass broader disciplines, including functions, charts, and scientific problems.
+Figure 1: (a) We showcase three examples of Text Redundancy (highlighted in red) within existing visual math benchmarks [9, 41, 63].
+(b) We report an ablation study by respectively removing the redundant texts and input diagrams on 120 randomly selected problems, for closed-sourced [47, 22, 3] and open-sourced [21, 38, 16] MLLMs.
 However, through our comprehensive observation and analysis, we identify three primary issues in current mathematical benchmarks for evaluating MLLMs:
@@
 Figure 2: Comparison of Visual Mathematical Reasoning by Three MLLMs.
-Despite the incorrect final answer, GPT-4V openai2023gpt4v , Gemini-Pro team2023gemini , and SPHINX-MoE gao2024sphinx exhibit different levels of quality in the intermediate reasoning process.
+Despite the incorrect final answer, GPT-4V [47], Gemini-Pro [22], and SPHINX-MoE [21] exhibit different levels of quality in the intermediate reasoning process.
 Therefore, in light of the issues discussed, we present MathVerse, a holistic and specialized visual math benchmark crafted to evaluate the multi-modal mathematical reasoning skills of MLLMs.
@@
 These delicately curated problems can indicate the various multi-modal capabilities of MLLMs, such as geometric element understanding, function curve perception, and numerical value recognition, which thoroughly unveils whether and how much they comprehend the visual diagram for mathematical reasoning.
-Second, to rigorously assess the visual Chain-of-Thought (CoT) capabilities wei2022chain , we propose a CoT Evaluation strategy for the step-by-step reasoning assessment of MLLMs.
+Second, to rigorously assess the visual Chain-of-Thought (CoT) capabilities [58], we propose a CoT Evaluation strategy for the step-by-step reasoning assessment of MLLMs.
 For each model’s output, we leverage GPT-4 to first extract several crucial steps exclusively from the solving process, deliberately omitting the input of the question and answer.
@@
 We present three examples in MathVerse for illustration.
-We conduct extensive experiments on MathVerse with popular closed-source openai2023gpt4v ; bai2023qwen ; team2023gemini and open-source liu2023improvedllava ; liu2024llavanext ; dong2024internlm ; gao2024sphinx MLLMs.
+We conduct extensive experiments on MathVerse with popular closed-source [47, 3, 22] and open-source [37, 38, 16, 21] MLLMs.
 Comparing different problem versions, we unveil that, most existing MLLMs struggle to understand math diagrams, relying heavily on textual questions.
-Therein, GPT-4V openai2023gpt4v achieves the best overall performance across different problem versions and subjects.
-Surprisingly, some of the MLLMs even attain much higher results without the diagram input, e.g., +5.1% for Qwen-VL-Max bai2023qwen and +5.6% for InternLM-XComposer2 dong2024internlm .
+Therein, GPT-4V [47] achieves the best overall performance across different problem versions and subjects.
+Surprisingly, some of the MLLMs even attain much higher results without the diagram input, e.g., +5.1% for Qwen-VL-Max [3] and +5.6% for InternLM-XComposer2 [16].
 With the fine-grained error analysis produced by our CoT evaluation strategy, we demonstrate such results are due to their deficient visual encoding capacity for mathematical diagrams, which instead acts as a distraction for problem-solving.
-In contrast, GPT-4V and ShareGPT4V Chen2023ShareGPT4VIL demonstrate relatively better comprehension of the visual content for mathematical reasoning.
+In contrast, GPT-4V and ShareGPT4V [12] demonstrate relatively better comprehension of the visual content for mathematical reasoning.
 Our experimental results suggest that inadequate mathematical visual interpretation capabilities represent the most significant impediment for MLLMs in addressing multi-modal math problems, indicating substantial potential for advancement.
@@
 Finally, in Section 2.3, we elaborate on the methodology of our proposed Chain-of-Thought (CoT) evaluation strategy.
-[…]
+1https://homework.study.com2https://www.ixl.com/math3https://mathspace.co/us
 2.1 Visual Math Dataset
@@
 Table 1: Key Statistics of MathVerse.
-Figure 4: Subject Distribution of MathVerse. Solid G: Solid Geometry, Plane G: Plane Geometry.
 Statistic Number Total questions 2,612 - Multiple-choice questions 1,631 (62.4%) - Free-form questions 981 (37.6%) - Newly collected questions 1,236 (47.3%) - Existing-dataset questions 1,376 (52.7%) - Questions with explanations 1,236 (47.3%) Total test samples 15,672 - Newly annotated samples 10,448 (66.7%) - Samples of each version 2,612 (16.7%) Number of unique images 2,420 (92.6%) Number of unique questions 2,573 (98.5%) Number of unique answers 847 (32.4%) Maximum question length 203 Maximum answer length 17 Average question length 35.7 Average answer length 1.4
@@
 The choice of these three subjects is not only due to their rigorous demands on multi-modal reasoning, but also for two other considerations.
-For one thing, as we specialize MathVerse in mathematical problem-solving, other peripheral tasks in MathVista Lu2023MathVistaEM are not included, e.g., statistical reasoning, table question-answering, and puzzle tests.
+For one thing, as we specialize MathVerse in mathematical problem-solving, other peripheral tasks in MathVista [41] are not included, e.g., statistical reasoning, table question-answering, and puzzle tests.
 For another, we expect the evaluation to fully display the reasoning capabilities of MLLMs with moderate-level mathematical knowledge.
 This avoids limiting their performance with overly complex domain-specific theorems or prior commonsense knowledge.
-Therefore, we deliberately focus the collected problems on the high school level, excluding advanced college-level disciplines like calculus and graph theory featured in MMMU yue2023mmmu .
+Therefore, we deliberately focus the collected problems on the high school level, excluding advanced college-level disciplines like calculus and graph theory featured in MMMU [63].
 Furthermore, expert annotators subdivide the problems into twelve fine-grained categories, as depicted in Figure 4, showcasing various dimensions of visual mathematical skills.
 Data Collection and Review Process.
-[…]
+Our collection procedure for high-quality visual math problems involves a rigorous selection from both pre-existing datasets and public question repositories.
+In the domain of plane geometry, we initially select 750 problems from GeoQA [9], 119 from GEOS [51], and 507 from Geometry3K [42], based on their original data quality and distribution.
+We exclude questions that are extremely simple or excessively complex, as well as those that appear dubious or lack necessary conditions.
+To enhance the diversity of question types and diagram styles, we further enrich our dataset with additional 370 plane geometry problems by manually collecting from other sources1,2,3.
+Given the scarcity of solid geometry and function-related problems in existing benchmarks, we purposefully gather these two types of problems (332 and 534, respectively) from new sources1,2,3 to address this gap.
+Problems that include multiple diagrams or require visual illustrations within solutions are excluded, considering the current limitations of MLLMs in resolving such information.
+Note that, all the newly collected problems (1,236) accompany detailed explanations.
+After the preliminary collection, we undertake a comprehensive review to verify the accuracy of the answers, ensure consistency between questions and diagrams, and confirm the relevance of each problem to the defined twelve categories.
+This meticulous review guarantees the dataset’s quality and precision.
 2.2 Whether MLLMs Truly See the Diagrams?
@@
 Figure 6: Examples of the CoT Evaluation Strategy for MathVerse.
-We present two outputs from Qwen-VL-Max bai2023qwen with our CoT evaluation strategy, which assesses the fine-grained reasoning capabilities with a detailed explanation for error analysis.
+We present two outputs from Qwen-VL-Max [3] with our CoT evaluation strategy, which assesses the fine-grained reasoning capabilities with a detailed explanation for error analysis.
 2.3 CoT Evaluation Strategy
@@
 Therefore, the binary ‘Correct’ or ‘Incorrect’ evaluative approach of existing benchmarks is inadequate to examine the depth and precision of the multi-step reasoning process.
-To this end, we propose a CoT evaluation strategy to thoroughly assess their mathematical CoT skills in visual contexts, involving two prompting phases with GPT-4(V) openai2023gpt4v ; OpenAI2023GPT4TR .
+To this end, we propose a CoT evaluation strategy to thoroughly assess their mathematical CoT skills in visual contexts, involving two prompting phases with GPT-4(V) [47, 46].
 Key-step Extraction.
@@
 We calculate the ‘All’ score without averaging the ‘Text Only’ version.
-‘CoT-E’ or ‘w/o’ denotes whether to employ the proposed CoT evaluation strategy or not.
+‘CoT-E’ or ‘Acc’ denotes whether to employ the proposed CoT evaluation strategy or not.
 The highest accuracy for closed-source and open-source MLLMs is marked in red and blue respectively.
-Model All Text Dominant Text Lite Text Only  Vision Intensive  Vision Dominant  Vision Only CoT-E w/o CoT-E w/o CoT-E w/o CoT-E w/o CoT-E w/o CoT-E w/o CoT-E w/o Baselines Random Chance - 12.4 - 12.4 - 12.4 - 12.4 - 12.4 - 12.4 - 12.4 Human - 64.9 - 71.2 - 70.9 - 41.7 - 61.4 - 68.3 - 66.7 LLMs ChatGPT ouyang2022training - - 51.3 33.3 38.5 18.9 51.3 33.3 - - - - - - GPT-4 OpenAI2023GPT4TR - - 63.4 46.5 40.7 20.7 63.4 46.5 - - - - - - Closed-source MLLMs Qwen-VL-Plus bai2023qwen 21.3 11.8 26.0 15.7 21.2 11.1 25.2 14.5 18.5 9.0 19.1 13.0 21.8 10.0 Gemini-Pro team2023gemini 35.3 22.3 39.8 27.6 34.7 23.7 44.5 27.9 32.0 19.4 36.8 20.3 33.3 20.5 Qwen-VL-Max bai2023qwen 37.2 24.8 42.8 30.3 37.7 24.8 47.9 32.2 33.6 20.6 35.9 23.3 35.9 25.1 GPT-4V openai2023gpt4v 54.4 38.3 63.1 52.1 56.6 40.9 60.3 46.1 51.4 34.9 50.8 33.6 50.3 29.8 Open-source MLLMs LLaMA-Adapter V2 gao2023llamaadapterv2 5.8 5.7 7.8 6.2 6.3 5.9 3.9 2.7 6.2 6.1 4.5 4.2 4.4 6.1 ImageBind-LLM han2023imagebind 10.0 9.3 13.2 11.4 11.6 11.3 12.9 11.7 9.8 8.9 11.8 11.2 3.5 3.4 mPLUG-Owl2 ye2023mplugowl2 10.3 4.6 11.6 6.6 11.4 6.3 13.8 6.1 11.1 6.3 9.4 5.6 8.0 4.9 MiniGPT-v2 chen2023minigpt 10.9 11.0 13.2 12.1 12.7 12.0 15.3 11.7 11.1 13.1 11.3 10.3 6.4 7.4 LLaVA-1.5 liu2023improvedllava 12.7 7.6 17.1 8.8 12.0 7.6 22.6 11.5 12.6 7.4 12.7 7.4 9.0 6.9 SPHINX-Plus gao2024sphinx 14.0 12.2 16.3 13.9 12.8 11.6 15.8 14.9 12.9 11.6 14.7 13.5 13.2 10.4 G-LLaVA gao2023g 15.7 16.6 22.2 20.9 20.4 20.7 21.6 21.1 16.5 17.2 12.7 14.6 6.6 9.4 LLaVA-NeXT liu2024llavanext 17.2 10.3 21.6 12.8 19.7 12.0 25.1 9.9 17.6 10.7 14.9 9.7 12.1 6.3 ShareGPT4V Chen2023ShareGPT4VIL 17.4 13.1 21.8 16.2 20.6 16.2 14.6 6.6 18.6 15.5 16.2 13.8 9.7 3.7 SPHINX-MoE gao2024sphinx 22.8 16.8 33.3 26.2 21.9 17.4 40.7 26.7 21.1 16.7 19.6 12.5 18.3 11.1 InternLM-XC2.
-dong2024internlm 25.9 16.3 36.9 20.2 28.3 14.3 42.5 24.5 20.1 14.2 24.4 17.5 19.8 15.2
+Model All Text Dominant Text Lite Text Only  Vision Intensive  Vision Dominant  Vision Only CoT-E Acc CoT-E Acc CoT-E Acc CoT-E Acc CoT-E Acc CoT-E Acc CoT-E Acc Baselines Random Chance - 12.4 - 12.4 - 12.4 - 12.4 - 12.4 - 12.4 - 12.4 Human - 64.9 - 71.2 - 70.9 - 41.7 - 61.4 - 68.3 - 66.7 LLMs ChatGPT [48] - - 51.3 33.3 38.5 18.9 51.3 33.3 - - - - - - GPT-4 [46] - - 63.4 46.5 40.7 20.7 63.4 46.5 - - - - - - Closed-source MLLMs Qwen-VL-Plus [3] 21.3 11.8 26.0 15.7 21.2 11.1 25.2 14.5 18.5 9.0 19.1 13.0 21.8 10.0 Gemini-Pro [22] 35.3 23.5 39.8 26.3 34.7 23.5 44.5 27.3 32.0 23.0 36.8 22.3 33.3 22.2 Qwen-VL-Max [3] 37.2 25.3 42.8 30.7 37.7 26.1 47.9 28.9 33.6 24.1 35.9 24.1 35.9 21.4 GPT-4V [47] 54.4 39.4 63.1 54.7 56.6 41.4 60.3 48.7 51.4 34.9 50.8 34.4 50.3 31.6 Open-source MLLMs LLaMA-Adapter V2 [20] 5.8 5.7 7.8 6.2 6.3 5.9 3.9 2.7 6.2 6.1 4.5 4.2 4.4 6.1 ImageBind-LLM [24] 10.0 9.2 13.2 11.4 11.6 11.3 12.9 11.7 9.8 8.9 11.8 11.2 3.5 3.4 mPLUG-Owl2 [62] 10.3 5.9 11.6 6.6 11.4 6.3 13.8 6.1 11.1 6.3 9.4 5.6 8.0 4.9 MiniGPT-v2 [11] 10.9 11.0 13.2 12.1 12.7 12.0 15.3 11.7 11.1 13.1 11.3 10.3 6.4 7.4 LLaVA-1.5 [37] 12.7 7.6 17.1 8.8 12.0 7.6 22.6 11.5 12.6 7.4 12.7 7.4 9.0 6.9 SPHINX-Plus [21] 14.0 12.2 16.3 13.9 12.8 11.6 15.8 14.9 12.9 11.6 14.7 13.5 13.2 10.4 G-LLaVA [19] 15.7 16.6 22.2 20.9 20.4 20.7 21.6 21.1 16.5 17.2 12.7 14.6 6.6 9.4 LLaVA-NeXT [38] 17.2 15.6 21.6 19.4 19.7 15.2 25.1 18.1 17.6 16.8 14.9 15.2 12.1 11.3 ShareGPT4V [12] 17.4 13.1 21.8 16.2 20.6 16.2 14.6 6.6 18.6 15.5 16.2 13.8 9.7 3.7 SPHINX-MoE [21] 22.8 15.0 33.3 22.2 21.9 16.4 40.7 18.3 21.1 14.8 19.6 12.6 18.3 9.1 InternLM-XC2.
+[16] 25.9 16.5 36.9 22.3 28.3 17.0 42.5 16.5 20.1 15.7 24.4 16.4 19.8 11.0
 3.1 Experimental Setup
@@
 Evaluation Models.
-We examine the performance of foundation models across three distinct categories on MathVerse: (a) Large Language Models (LLMs) as the text-only baseline, which only take textual questions as input, including ChatGPT OpenAI2023ChatGPT and GPT-4 OpenAI2023GPT4TR , (b) Closed-source MLLMs, represented by models like GPT-4V openai2023gpt4v , Gemini-Pro team2023gemini , Qwen-VL-Max bai2023qwen , and Qwen-VL-Plus, and (c) Open-source MLLMs, featuring models such as LLaVA-1.5 liu2023improvedllava (Vicuna-13B vicuna2023 ), LLaVA-NeXT liu2024llavanext (Vicuna-13B), SPHINX-MoE gao2024sphinx (Mixtral-8×7B albert24mixtral ), SPHINX-Plus (LLaMA2-13B touvron2023llama2 ), InternLM-XComposer2 dong2024internlm (InternLM2-7B team2023internlm ), LLaMA-Adapter V2 gao2023llamaadapterv2 (LLaMA-7B touvron2023llama ), ImageBind-LLM han2023imagebind (LLaMA-7B), MiniGPT-v2 chen2023minigpt (LLaMA2-7B), mPLUG-Owl2 ye2023mplugowl2 (LLaMA-7B), G-LLaVA gao2023g (LLaMA2-7B), and ShareGPT-4V Chen2023ShareGPT4VIL (Vicuna-13B).
+We examine the performance of foundation models across three distinct categories on MathVerse: (a) Large Language Models (LLMs) as the text-only baseline, which only take textual questions as input, including ChatGPT [45] and GPT-4 [46], (b) Closed-source MLLMs, represented by models like GPT-4V [47], Gemini-Pro [22], Qwen-VL-Max [3], and Qwen-VL-Plus, and (c) Open-source MLLMs, featuring models such as LLaVA-1.5 [37] (Vicuna-13B [13]), LLaVA-NeXT [38] (Vicuna-13B), SPHINX-MoE [21] (Mixtral-8×7B [28]), SPHINX-Plus (LLaMA2-13B [56]), InternLM-XComposer2 [16] (InternLM2-7B [54]), LLaMA-Adapter V2 [20] (LLaMA-7B [55]), ImageBind-LLM [24] (LLaMA-7B), MiniGPT-v2 [11] (LLaMA2-7B), mPLUG-Owl2 [62] (LLaMA-7B), G-LLaVA [19] (LLaMA2-7B), and ShareGPT-4V [12] (Vicuna-13B).
 Table 3: Mathematical Evaluation on Different Subjects and Subfields in MathVerse’s testmini Set.
@@
 The highest accuracy for closed-source and open-source MLLMs is marked in red and blue respectively.
-Model All Plane Geometry Solid Geometry Functions All Len Area Angle Anal Apply All Len Area Vol All Coord Prop Exp Apply Closed-source MLLMs Qwen-VL-Plus bai2023qwen 21.3 17.3 19.1 16.4 16.1 23.6 13.2 24.8 18.1 18.7 33.4 31.3 52.5 25.1 10.8 50.3 Gemini-Pro team2023gemini 35.3 33.0 32.2 42.6 28.4 30.2 32.3 33.4 35.0 29.3 36.1 28.3 25.7 26.6 10.8 51.3 Qwen-VL-Max bai2023qwen 37.2 38.4 41.7 46.4 32.6 40.6 38.7 33.7 25.4 28.3 42.6 38.4 43.7 35.5 13.6 61.0 GPT-4V openai2023gpt4v 54.4 56.9 60.8 63.4 52.6 48.5 60.9 50.2 54.8 39.9 56.8 52.8 72.3 47.1 30.9 70.1 Open-source MLLMs LLaMA-Adapter V2  gao2023llamaadapterv2 5.8 5.9 4.0 5.9 6.6 13.4 3.3 4.6 5.3 3.1 5.7 6.2 6.7 6.1 4.5 7.9 ImageBind-LLM han2023imagebind 10.0 9.7 12.1 9.9 9.2 10.2 4.8 4.6 4.9 3.5 5.3 14.9 12.3 13.8 4.6 25.9 mPLUG-Owl2 ye2023mplugowl2 10.3 7.7 8.2 6.0 5.7 12.4 10.6 11.0 9.2 6.7 15.7 17.4 22.8 18.6 5.3 22.2 MiniGPT-v2 chen2023minigpt 10.9 11.6 10.0 9.8 14.3 9.1 11.8 1.7 2.2 1.6 0.5 11.2 4.2 15.7 4.0 21.1 LLaVA-1.5 liu2023improvedllava 12.7 11.8 13.1 15.1 9.7 9.4 13.2 10.6 12.1 8.7 11.6 14.8 18.8 12.7 9.5 23.7 SPHINX-Plus gao2024sphinx 14.0 14.4 14.2 10.5 14.1 16.5 16.8 7.0 7.2 6.1 7.6 17.9 11.1 19.1 6.3 27.7 G-LLaVA gao2023g 15.7 20.2 17.3 13.6 26.5 5.9 23.1 5.0 10.3 4.4 3.1 9.2 9.1 9.1 1.3 15.5 LLaVA-NeXT liu2024llavanext 17.2 15.9 14.8 13.1 16.3 17.7 17.8 19.6 33.3 11.7 12.6 23.1 24.5 23.4 8.0 33.1 ShareGPT4V Chen2023ShareGPT4VIL 17.4 16.9 16.2 17.9 16.9 12.2 21.1 15.0 13.6 10.9 19.7 20.2 19.9 22.2 8.4 25.8 SPHINX-MoE gao2024sphinx 22.8 24.5 26.3 28.4 21.1 26.6 24.4 15.8 9.4 10.7 26.3 19.5 23.5 19.3 9.2 30.3 InternLM-XC2.
-dong2024internlm 25.9 26.2 27.1 29.7 20.6 18.5 22.2 20.1 34.5 14.1 25.2 23.7 24.4 24.9 10.6 36.3
+Model All Plane Geometry Solid Geometry Functions All Len Area Angle Anal Apply All Len Area Vol All Coord Prop Exp Apply Closed-source MLLMs Qwen-VL-Plus [3] 21.3 17.3 19.1 16.4 16.1 23.6 13.2 24.8 18.1 18.7 33.4 31.3 52.5 25.1 10.8 50.3 Gemini-Pro [22] 35.3 33.0 32.2 42.6 28.4 30.2 32.3 33.4 35.0 29.3 36.1 28.3 25.7 26.6 10.8 51.3 Qwen-VL-Max [3] 37.2 38.4 41.7 46.4 32.6 40.6 38.7 33.7 25.4 28.3 42.6 38.4 43.7 35.5 13.6 61.0 GPT-4V [47] 54.4 56.9 60.8 63.4 52.6 48.5 60.9 50.2 54.8 39.9 56.8 52.8 72.3 47.1 30.9 70.1 Open-source MLLMs LLaMA-Adapter V2  [20] 5.8 5.9 4.0 5.9 6.6 13.4 3.3 4.6 5.3 3.1 5.7 6.2 6.7 6.1 4.5 7.9 ImageBind-LLM [24] 10.0 9.7 12.1 9.9 9.2 10.2 4.8 4.6 4.9 3.5 5.3 14.9 12.3 13.8 4.6 25.9 mPLUG-Owl2 [62] 10.3 7.7 8.2 6.0 5.7 12.4 10.6 11.0 9.2 6.7 15.7 17.4 22.8 18.6 5.3 22.2 MiniGPT-v2 [11] 10.9 11.6 10.0 9.8 14.3 9.1 11.8 1.7 2.2 1.6 0.5 11.2 4.2 15.7 4.0 21.1 LLaVA-1.5 [37] 12.7 11.8 13.1 15.1 9.7 9.4 13.2 10.6 12.1 8.7 11.6 14.8 18.8 12.7 9.5 23.7 SPHINX-Plus [21] 14.0 14.4 14.2 10.5 14.1 16.5 16.8 7.0 7.2 6.1 7.6 17.9 11.1 19.1 6.3 27.7 G-LLaVA [19] 15.7 20.2 17.3 13.6 26.5 5.9 23.1 5.0 10.3 4.4 3.1 9.2 9.1 9.1 1.3 15.5 LLaVA-NeXT [38] 17.2 15.9 14.8 13.1 16.3 17.7 17.8 19.6 33.3 11.7 12.6 23.1 24.5 23.4 8.0 33.1 ShareGPT4V [12] 17.4 16.9 16.2 17.9 16.9 12.2 21.1 15.0 13.6 10.9 19.7 20.2 19.9 22.2 8.4 25.8 SPHINX-MoE [21] 22.8 24.5 26.3 28.4 21.1 26.6 24.4 15.8 9.4 10.7 26.3 19.5 23.5 19.3 9.2 30.3 InternLM-XC2.
+[16] 25.9 26.2 27.1 29.7 20.6 18.5 22.2 20.1 34.5 14.1 25.2 23.7 24.4 24.9 10.6 36.3
 Implementation Details.
 All our experiments are conducted under a zero-shot setting, showcasing the generalization capacity of MLLMs for mathematical reasoning, without few-shot prompting or further fine-tuning.
-By default, we employ the Chain-of-Thought (CoT) prompting technique wei2022chain , which encourages MLLMs to perform complete reasoning steps for a fine-grained evaluation.
+By default, we employ the Chain-of-Thought (CoT) prompting technique [58], which encourages MLLMs to perform complete reasoning steps for a fine-grained evaluation.
 A baseline representing random chance is established for comparison, for which we select one option at random for multiple-choice questions and utilize empty for free-form questions.
@@
 As the text-only LLMs can only take text questions as input, we evaluate them with the first three problem versions, i.e., Text Dominant, Text Lite, and Text Only.
-For the ‘w/o’ results, we utilize the template in MathVista Lu2023MathVistaEM to prompt GPT-4 OpenAI2023GPT4TR for answer extraction, and directly score the final answer without the intermediate reasoning process.
+For the ‘w/o’ results, we utilize the template in MathVista [41] to prompt GPT-4 [46] for answer extraction, and directly score the final answer without the intermediate reasoning process.
 3.2 Experimental Analysis
@@
 Figure 7: A Typical Visual Perception Error by our CoT Evaluation Strategy.
-The example is an output from Gemini-Pro team2023gemini , where the correct reasoning of the Text-only version is distracted by the visual perception error within the diagram.
+The example is an output from Gemini-Pro [22], where the correct reasoning of the Text-only version is distracted by the visual perception error within the diagram.
 MLLMs are Moderately Effective at Perceiving IP.
@@
 Mathematical Training Benefits the Performance.
-In addition to foundational visual instruction-following datasets, both SPHINX-MoE and InternLM-XComposer2 extend their training regimes to include specialized mathematical problems that are either text-only or visual, such as MathQA amini2019mathqa , Geometry3K Lu2021InterGPSIG , and MathInstruct yue2023mammoth .
+In addition to foundational visual instruction-following datasets, both SPHINX-MoE and InternLM-XComposer2 extend their training regimes to include specialized mathematical problems that are either text-only or visual, such as MathQA [2], Geometry3K [43], and MathInstruct [64].
 This approach of math-specific tuning contributes to their leading performance in MathVerse.
@@
 3.3 Error Analysis
-Figure 9: Distribution of GPT-4V’s openai2023gpt4v Errors in Reasoning and Answers.
+Figure 9: Distribution of GPT-4V’s [47] Errors in Reasoning and Answers.
 For the six problem versions in MathVerse, we provide the statistics of errors made by GPT-4V based on their occurrence in answers (‘Ans.’) and reasoning processes (‘Rea.’).
-To delve into the fine-grained predictions, we select the best-performing MLLM, GPT-4V openai2023gpt4v , to understand its modes of success and failure.
+To delve into the fine-grained predictions, we select the best-performing MLLM, GPT-4V [47], to understand its modes of success and failure.
 Our proposed CoT evaluation strategy has produced a detailed assessment of model output, including step-wise scores and explanation, reducing extensive manual effort in identifying and analyzing errors.
@@
 As we remove the content within textual questions and enrich the visual diagram, e.g., from Text Dominant and Lite to Vision Dominant and Only, we observe a progressive increase in the error rate of ‘incorrect final answer with incorrect CoT reasoning’, indicating that MLLMs are challenged to conduct high-quality intermediate reasoning by capturing more information from the visual input.
-Figure 10: Distribution of GPT-4V’s openai2023gpt4v Errors within Different Types.
+Figure 10: Distribution of GPT-4V’s [47] Errors within Different Types.
 We present the statistics of four error types by GPT-4V in the six problem versions, i.e., Visual Perception Error, Reasoning Error, Calculation Error, and Knowledge Error.
@@
 Multi-modal Large Language Models (MLLMs),
-building upon the prevalence of Large Language Models (LLMs) touvron2023llama ; touvron2023llama2 ; OpenAI2023ChatGPT ; albert24mixtral ; brown2020language and large vision models Radford2021LearningTV ; kirillov2023segment ; zhang2023personalize ; zhang2023prompt ; zhang2023learning , have become increasingly prominent in the field.
-They extend LLMs to tackle a diverse range of tasks and domains, including the mainstream 2D images li2022blip ; instructblip ; alayrac2022flamingo ; li2023mimic and other modalities, such as 3D point clouds guo2023point ; xu2023pointllm ; hong20243d , audio han2023imagebind ; su2023pandagpt , and video zhang2023video ; chen2023videollm .
-Noteworthy examples like OpenAI’s GPT-4V openai2023gpt4v and Google’s Gemini team2023gemini exhibit exceptional visual understanding and reasoning capabilities, setting new benchmarks in multi-modal performance.
+building upon the prevalence of Large Language Models (LLMs) [55, 56, 45, 28, 4] and large vision models [49, 29, 68, 67, 69], have become increasingly prominent in the field.
+They extend LLMs to tackle a diverse range of tasks and domains, including the mainstream 2D images [35, 15, 1, 31] and other modalities, such as 3D point clouds [23, 60, 27], audio [24, 52], and video [65, 6].
+Noteworthy examples like OpenAI’s GPT-4V [47] and Google’s Gemini [22] exhibit exceptional visual understanding and reasoning capabilities, setting new benchmarks in multi-modal performance.
 However, their closed-source nature poses a barrier to the broader application and development of MLLMs.
 Concurrently, another line of work is dedicated to exploring advanced MLLMs open-source to the community.
-Prior efforts like LLaMA-Adapter zhang2024llamaadapter ; gao2023llamaadapterv2 , LLaVA liu2023llava ; liu2024llavanext ; liu2023improvedllava , and MiniGPT-4 zhu2023minigpt ; chen2023minigpt leverage a frozen CLIP Radford2021LearningTV model for image encoding, and inject the visual cues into LLaMA touvron2023llama for multi-modal instruction tuning.
-The subsequent mPLUG-Owl ye2023mplugowl ; ye2023mplugowl2 , Qwen-VL bai2023qwen , InternLM-XComposer dong2024internlm , and SPHINX lin2023sphinx ; gao2024sphinx further push the frontier of MLLMs in understanding and generalizing across visual contexts.
-Despite comprehensive benchmarks fu2023mme ; liu2023mmbench ; Li2023SEEDBenchBM ; xu2023lvlm on general visual instruction-following scenarios, the specific potential of MLLMs for visual mathematical problem-solving remains under-explored.
+Prior efforts like LLaMA-Adapter [66, 20], LLaVA [39, 38, 37], and MiniGPT-4 [72, 11] leverage a frozen CLIP [49] model for image encoding, and inject the visual cues into LLaMA [55] for multi-modal instruction tuning.
+The subsequent mPLUG-Owl [61, 62], Qwen-VL [3], InternLM-XComposer [16], and SPHINX [36, 21] further push the frontier of MLLMs in understanding and generalizing across visual contexts.
+Despite comprehensive benchmarks [17, 40, 33, 59] on general visual instruction-following scenarios, the specific potential of MLLMs for visual mathematical problem-solving remains under-explored.
 In this paper, we introduce the MathVerse benchmark to comprehensively evaluate the visual mathematical reasoning and diagram understanding skills of MLLMs, providing unique perspectives for future research directions.
@@
 have emerged as a significant area of focus, posing considerable challenges for large foundational models, e.g., LLMs and MLLMs.
-Initially, datasets in this realm are designed to address basic algebraic hendrycksmath2021 and arithmetic Roy2016SolvingGA word problems, which are relatively limited in scope and volume.
-Subsequent efforts, including MATH hendrycksmath2021 , GSM8K cobbe2021training , and MMLU hendryckstest2021 , expand the range and quality of textual mathematical problems.
-These datasets feature a broader spectrum of difficulties, establishing a robust benchmark for the evaluation of general and math-specific LLMs zhou2023solving ; yue2023mammoth ; wang2024mathcoder ; gao2023g ; luo2023wizardmath .
+Initially, datasets in this realm are designed to address basic algebraic [26] and arithmetic [50] word problems, which are relatively limited in scope and volume.
+Subsequent efforts, including MATH [26], GSM8K [14], and MMLU [25], expand the range and quality of textual mathematical problems.
+These datasets feature a broader spectrum of difficulties, establishing a robust benchmark for the evaluation of general and math-specific LLMs [71, 64, 57, 19, 44].
 Besides the text-only assessment, there is a growing demand for comparable, high-quality benchmarks for evaluating mathematical problem-solving in visual contexts, with the rapid progress of MLLMs.
-There are prior attempts, such as GeoQA Chen2021GeoQAAG , UniGeo Chen2022UniGeoUG , and Geometry3K Lu2021InterGPSIG , which focused exclusively on geometric problems.
-The recently proposed MathVista Lu2023MathVistaEM broadens the scope to incorporate a variety of multi-modal tasks involving mathematical reasoning, and MMMU yue2023mmmu covers college-level questions demanding intricate, domain-specific knowledge.
+There are prior attempts, such as GeoQA [10], UniGeo [8], and Geometry3K [43], which focused exclusively on geometric problems.
+The recently proposed MathVista [41] broadens the scope to incorporate a variety of multi-modal tasks involving mathematical reasoning, and MMMU [63] covers college-level questions demanding intricate, domain-specific knowledge.
 However, our analysis identifies three main shortcomings within the current visual math benchmarks, as elaborated in Section 1 of the main paper.
@@
 Table 4: The Release Time and Model Source of MLLMs Used in MathVerse.
-Model Release Time
-Source
-ChatGPT ouyang2022training 2022-11
-https://platform.openai.com/
-GPT-4 OpenAI2023GPT4TR 2023-03
-https://platform.openai.com/
-Qwen-VL-Plus bai2023qwen 2023-11
-https://help.aliyun.com/zh/dashscope/developer-reference/vl-plus-quick-start
-Gemini-Pro team2023gemini 2023-12
-https://ai.google.dev/
-Qwen-VL-Max bai2023qwen 2024-01
-https://help.aliyun.com/zh/dashscope/developer-reference/vl-plus-quick-start
-GPT-4V openai2023gpt4v 2023-09
-https://platform.openai.com/
-LLaMA-Adapter V2 gao2023llamaadapterv2 2023-04
-https://github.com/OpenGVLab/LLaMA-Adapter/tree/main/llama_adapter_v2_multimodal7b
-LLaVA-1.5 liu2023improvedllava 2023-10
-https://huggingface.co/liuhaotian/llava-v1.5-13b
-MiniGPT-v2 chen2023minigpt 2023-10
-https://github.com/Vision-CAIR/MiniGPT-4
-mPLUG-Owl2 ye2023mplugowl2 2023-11
-https://huggingface.co/MAGAer13/mplug-owl2-llama2-7b
-G-LLaVA gao2023g 2023-12
-https://github.com/pipilurj/G-LLaVA/tree/main
-ImageBind-LLM han2023imagebind 2023-05
-https://github.com/OpenGVLab/LLaMA-Adapter/tree/main/imagebind_LLM
-ShareGPT4V Chen2023ShareGPT4VIL 2023-11
-https://huggingface.co/Lin-Chen/ShareGPT4V-13B
-SPHINX-Plus lin2023sphinx 2023-11
-https://huggingface.co/Alpha-VLLM/LLaMA2-Accessory/tree/main/finetune/mm/SPHINX/SPHINX-v2-1k
-LLaVA-NeXT liu2024llavanext 2024-01
-https://huggingface.co/liuhaotian/llava-v1.6-vicuna-13b
-SPHINX-MoE gao2024sphinx 2024-01
-https://huggingface.co/Alpha-VLLM/LLaMA2-Accessory/tree/main/finetune/mm/SPHINX/SPHINX-MoE
-InternLM-XComposer2 dong2024internlm 2024-01
-https://huggingface.co/internlm/internlm-xcomposer2-vl-7b
 Appendix B Additional Experimental Details
@@
 Question Prompt
-Free-form Question
-Please first conduct reasoning, and then answer the question and provide the final value, e.g., 1, 2.5, 300, at the end.
+Free-form Question Please first conduct reasoning, and then answer the question and provide the final value, e.g., 1, 2.5, 300, at the end.
 Question: $\{question\}$
-Multiple-choice Question
-Please first conduct reasoning, and then answer the question and provide the correct option letter, e.g., A, B, C, D, at the end.
+Multiple-choice Question Please first conduct reasoning, and then answer the question and provide the correct option letter, e.g., A, B, C, D, at the end.
 Question: $\{question\}$
@@
 Question Prompt
-Free-form Question
-According to the question shown in the image, please first conduct reasoning, and then answer the question and provide the final value, e.g., 1, 2.5, 300, at the end.
-Multiple-choice Question
-According to the question shown in the image, please first conduct reasoning, and then answer the question and provide the correct option letter, e.g., A, B, C, D, at the end.
+Free-form Question According to the question shown in the image, please first conduct reasoning, and then answer the question and provide the final value, e.g., 1, 2.5, 300, at the end.
+Multiple-choice Question According to the question shown in the image, please first conduct reasoning, and then answer the question and provide the correct option letter, e.g., A, B, C, D, at the end.
 Prompt for Response Generation.
@@
 Figure 11: Manual Annotations for Function Problems in MathVerse.
-We provide detailed annotations, e.g., function expression and properties, for the diagrams of 534 function problems, which benefits the accuracy of GPT-4V openai2023gpt4v for CoT evaluation.
+We provide detailed annotations, e.g., function expression and properties, for the diagrams of 534 function problems, which benefits the accuracy of GPT-4V [47] for CoT evaluation.
 Prompt for the CoT Evaluation.
-Our proposed CoT evaluation contains two steps, i.e., key-step extraction and multi-step scoring, which prompt GPT-4 OpenAI2023GPT4TR and GPT-4V openai2023gpt4v , respectively.
+Our proposed CoT evaluation contains two steps, i.e., key-step extraction and multi-step scoring, which prompt GPT-4 [46] and GPT-4V [47], respectively.
 The input configuration is listed in Table 7.
@@
 Table 7: Configuration for the CoT Evaluation Strategy.
-We conduct two evaluation phases, respectively by prompting the text-only GPT-4 OpenAI2023GPT4TR and GPT-4V openai2023gpt4v .
+We conduct two evaluation phases, respectively by prompting the text-only GPT-4 [46] and GPT-4V [47].
 The symbol ‘XXX’ denotes the given one-shot sample, abbreviated for brevity.
@@
 Phase Input Prompt
-Key-step Extraction (GPT-4) Model Output
-I will give you a detailed solving procedure or a single answer for a math problem.
-If it is a procedure, you need to extract the key solution steps and list them accordingly in markdown syntax. If it is just a single answer, output the answer directly.
-Here are examples:
-– Model output: XXX
-– Extracted: 1. XXX 2. XXX 3. XXX
-– Model output: 2.2
-– Extracted: The single answer is 2.2
-Here is what you need to extract:
-Model output: $\{model\ output\}$
-Extracted:
-Multi-step Scoring (GPT-4V) Extracted Steps Question Diagram Answer Annotation
-I will first give you a visual math problem, including the question, diagram, ground-truth answer, and detailed annotation of the diagram, and then give you a model output containing multiple key solution steps.
-Please think step by step and output the Average score, along with the Final answer score in the end, as described below:
-– Average score: Evaluate, based on the given question, answer, diagram, and diagram annotation, whether each solution step is correct in logical reasoning, visual perception, and numerical computation, with an incorrect score of 0 and a correct score of 1.
+Key-step Extraction (GPT-4) Model Output I will give you a detailed solving procedure or a single answer for a math problem.
+If it is a procedure, you need to extract the key solution steps and list them accordingly in markdown syntax.
+If it is just a single answer, output the answer directly.
+Here are examples: – Model output: XXX – Extracted: 1.
+XXX 2.
+XXX 3.
+XXX – Model output: 2.2 – Extracted: The single answer is 2.2 Here is what you need to extract: Model output: $\{model\ output\}$ Extracted:
+Multi-step Scoring (GPT-4V) Extracted Steps Question Diagram Answer Annotation I will first give you a visual math problem, including the question, diagram, ground-truth answer, and detailed annotation of the diagram, and then give you a model output containing multiple key solution steps.
+Please think step by step and output the Average score, along with the Final answer score in the end, as described below: – Average score: Evaluate, based on the given question, answer, diagram, and diagram annotation, whether each solution step is correct in logical reasoning, visual perception, and numerical computation, with an incorrect score of 0 and a correct score of 1.
 Then, calculate the average score of multiple steps.
@@
 – If the model output only includes a single step or answer, the Average score and Final answer score are the same.
-Question: $\{question\}$
-Ground-truth answer: $\{answer\}$
-Diagram annotation: $\{annotation\}$
-Model output: $\{extracted\ steps\}$
-Average score:
-Final answer score:
+Question: $\{question\}$ Ground-truth answer: $\{answer\}$ Diagram annotation: $\{annotation\}$ Model output: $\{extracted\ steps\}$ Average score: Final answer score:
 Human Performance Assessment.
@@
 Therefore, we do not report the CoT evaluation results for human performance, alongside the ‘Random Chance’ baseline.
-[…]
+1https://homework.study.com2https://www.ixl.com/math3https://mathspace.co/us4https://mathpix.com
 Appendix C More Dataset Details
@@
 Data Collection.
-[…]
+We comprehensively collect visual math problems from existing datasets [43, 9, 51] and public question repositories1,2,3.
+We specifically select high-quality plane geometric problems from current benchmarks, which showcase various question types, moderate question length, diverse diagram styles, and appropriate solving difficulty.
+For the manually collected problems of three subjects (plane geometry, solid geometry, and functions), we apply the Mathpix tool4 to accurately extract the question texts, diagrams, explanations, and answers from the website.
+We strictly comply with copyright and licensing rules, ensuring that we refrain from using data from sites that forbid copying and redistribution.
+After the initial collection, we obtain around 3.5K visual math problems, with 1.5K from existing datasets and 2K newly collected.
 Data Categorization and Review.
@@
 Table 8: Statistics of MathVerse.
+Statistic Number Total questions 2,612 - Subjects/subfields 3/12 - Multiple-choice questions 1,631 (62.4%) - Free-form questions 981 (37.6%) - Newly collected questions 1,236 (47.3%) - Existing-dataset questions 1,376 (52.7%) - Questions with explanations 1,236 (47.3%) - Newly annotated questions 534 (20.4%) Multiple-choice question - Proportion of answer A 585 (22.4%) - Proportion of answer B 828 (31.7%) - Proportion of answer C 703 (26.9%) - Proportion of answer D 444 (17.0%) - Proportion of answer E&F 52 (2.0%) Total test samples 15,672 - Newly annotated samples 10,448 (66.7%) - Newly annotated diagrams 5,224 (33.3%) - Samples of each version 2,612 (16.7%) Number of unique images 2,420 (92.6%) Number of unique questions 2,573 (98.5%) Number of unique answers 847 (32.4%)
 Table 9: Length of Different Problem Versions in MathVerse.
-Statistic Number Total questions 2,612  - Subjects/subfields 3/12  - Multiple-choice questions 1,631 (62.4%)  - Free-form questions 981 (37.6%)  - Newly collected questions 1,236 (47.3%)  - Existing-dataset questions 1,376 (52.7%)  - Questions with explanations 1,236 (47.3%)  - Newly annotated questions 534 (20.4%) Multiple-choice question  - Proportion of answer A 585 (22.4%)  - Proportion of answer B 828 (31.7%)  - Proportion of answer C 703 (26.9%)  - Proportion of answer D 444 (17.0%)  - Proportion of answer E&F 52 (2.0%) Total test samples 15,672  - Newly annotated samples 10,448 (66.7%)  - Newly annotated diagrams 5,224 (33.3%)  - Samples of each version 2,612 (16.7%) Number of unique images 2,420 (92.6%) Number of unique questions 2,573 (98.5%) Number of unique answers 847 (32.4%)
 Problem Version    Word    Character Text Dominant & Text Only - Maximum question length 203 1,311 - Maximum answer length 17 102 - Average question length 35.7 204.8 - Average answer length 1.4 6.3 Text Lite - Maximum question length 179 1,173 - Maximum answer length 17 102 - Average question length 22 133.8 - Average answer length 1.4 6.3 Vision Intensive - Maximum question length 171 1,126 - Maximum answer length 17 102 - Average question length 18.8 116.8 - Average answer length 1.4 6.3 Vision Dominant - Maximum question length 176 1,132 - Maximum answer length 17 102 - Average question length 17.6 123.5 - Average answer length 1.4 6.3
-Table 9: Length of Different Problem Versions in MathVerse.
 Problem Length Variance.
@@
 Appendix D Comparison to Current Benchmarks
-In this section, we offer a detailed comparison between MathVerse and existing multi-modal mathematical benchmarks, i.e., geometry-specific benchmarks chen2021geoqa ; cao2022augmented ; chen2022unigeo ; lu2021inter ; seo2015solving , MathVista Lu2023MathVistaEM , and MMMU yue2023mmmu , from the following four aspects:
+In this section, we offer a detailed comparison between MathVerse and existing multi-modal mathematical benchmarks, i.e., geometry-specific benchmarks [9, 5, 7, 42, 51], MathVista [41], and MMMU [63], from the following four aspects:
 The Investigation of Diagram Interpretation Capacity.
@@
 By comparing the performance variance between different problem versions, we can thoroughly investigate the mathematical diagram interpretation capabilities of MLLMs for the first time.
-Figure 18: Diagram Examples of Math-related Tasks in MathVista Lu2023MathVistaEM .
+Figure 18: Diagram Examples of Math-related Tasks in MathVista [41].
 These tasks are not strongly correlated to the mathematical reasoning skills of MLLMs, probably skewing the assessment emphasis towards visual math problems.
-Figure 19: Diagram Examples with Required Knowledge in MMMU yue2023mmmu .
+Figure 19: Diagram Examples with Required Knowledge in MMMU [63].
 These math problems demand MLLMs to comprehend college-level domain knowledge, potentially hindering them from fully exerting mathematical reasoning skills.
@@
 Table 10: Number of Test Samples in Different Benchmarks.
-Benchmark
-GEOS
-Geo3K
-GeoQA+
-MathVista
-MMMU-Math
-MathVerse
-Test Samples
-119
-601
-755
-6,141
-540
-15,672
+Benchmark GEOS Geo3K GeoQA+ MathVista MMMU-Math MathVerse
+Test Samples 119 601 755 6,141 540 15,672
 Appendix E Limitation and Future Work
@@
 These categorization approaches evaluate the capabilities of MLLMs from multiple dimensions.
-Nevertheless, it is also meaningful to further divide the problems based on their difficulty levels, akin to MATH hendrycksmath2021 , a text-only benchmark defining five levels of difficulty.
+Nevertheless, it is also meaningful to further divide the problems based on their difficulty levels, akin to MATH [26], a text-only benchmark defining five levels of difficulty.
 This additional layer of differentiation can provide deeper insights into the problem-solving abilities of MLLMs across a spectrum of challenges, which we leave as future work.
@@
 Moreover, the problems in MathVerse and other current mathematical benchmarks are mainly in English.
-Given that some multilingual MLLMs han2023imagebind ; bai2023qwen have been developed, existing evaluation cannot reveal their full capabilities when confined to a single language.
+Given that some multilingual MLLMs [24, 3] have been developed, existing evaluation cannot reveal their full capabilities when confined to a single language.
 The incorporation of multilingual visual math problems would not only extend the dataset’s global applicability, but also enhance the assessment of MLLMs for linguistic diversity and understanding.
@@
 In Section F.1, we showcase the meticulously transformed six versions of visual math problems.
-In Section F.2, we compare the response of different MLLMs on Text-lite problems, including GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx .
+In Section F.2, we compare the response of different MLLMs on Text-lite problems, including GPT-4V [47], LLaVA-NeXT [38], and SPHINX-MoE [21].
 Specifically, we present the key-step extraction output by the CoT evaluation, and mark the multi-step scoring results aside.
@@
 Please refer to Figures 16$\sim$21.
-Figure 23: Response Comparison of GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx .
+Figure 23: Response Comparison of GPT-4V [47], LLaVA-NeXT [38], and SPHINX-MoE [21].
 We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
-Figure 24: Response Comparison of GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx .
+Figure 24: Response Comparison of GPT-4V [47], LLaVA-NeXT [38], and SPHINX-MoE [21].
 We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
-Figure 25: Response Comparison of GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx .
+Figure 25: Response Comparison of GPT-4V [47], LLaVA-NeXT [38], and SPHINX-MoE [21].
 We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
-Figure 26: Response Comparison of GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx .
+Figure 26: Response Comparison of GPT-4V [47], LLaVA-NeXT [38], and SPHINX-MoE [21].
 We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
-Figure 27: Response Comparison of GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx .
+Figure 27: Response Comparison of GPT-4V [47], LLaVA-NeXT [38], and SPHINX-MoE [21].
 We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
-Figure 28: Response Comparison of GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx .
+Figure 28: Response Comparison of GPT-4V [47], LLaVA-NeXT [38], and SPHINX-MoE [21].
 We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
@@
 Please refer to Figures 22$\sim$27.
-Figure 29: Response of Different Problem Versions by GPT-4V openai2023gpt4v .
+Figure 29: Response of Different Problem Versions by GPT-4V [47].
 By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams.
 We only showcase the Text-dominant version of the problem for brevity.
-Figure 30: Response of Different Problem Versions by GPT-4V openai2023gpt4v .
+Figure 30: Response of Different Problem Versions by GPT-4V [47].
 By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams.
 We only showcase the Text-dominant version of the problem for brevity.
-Figure 31: Response of Different Problem Versions by GPT-4V openai2023gpt4v .
+Figure 31: Response of Different Problem Versions by GPT-4V [47].
 By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams.
 We only showcase the Text-dominant version of the problem for brevity.
-Figure 32: Response of Different Problem Versions by GPT-4V openai2023gpt4v .
+Figure 32: Response of Different Problem Versions by GPT-4V [47].
 By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams.
 We only showcase the Text-dominant version of the problem for brevity.
-Figure 33: Response of Different Problem Versions by GPT-4V openai2023gpt4v .
+Figure 33: Response of Different Problem Versions by GPT-4V [47].
 By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams.
 We only showcase the Text-dominant version of the problem for brevity.
-Figure 34: Response of Different Problem Versions by GPT-4V openai2023gpt4v .
+Figure 34: Response of Different Problem Versions by GPT-4V [47].
 By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams.